### v1

# Uniform approximation on certain polynomial polyhedra in $\mathbb{C}^{2}$

Abstract.
In this paper we extend the dichotomy given by Samuelsson and Wold [28] that can be thought of as an analogue of the Wermer maximality theorem in $\mathbb{C}^{2}$ for certain polynomial polyhedra. We consider complex non-degenerate simply connected polynomial polyhedra of the form $\Omega:=\{z\in\mathbb{C}^{2}:|p_{1}(z)|<1,|p_{2}(z)|<1\}$ such that $\overline{\Omega}$ is compact. Under a mild condition of the polynomials $p_{1}$ and $p_{2}$, we prove that either the uniform algebra, generated by polynomials and some continuous functions $f_{1},\dots,f_{N}$ on the distinguished boundary that extends as pluriharmonic functions on $\Omega$, is all continuous functions on the distinguished boundary or there exists an algebraic variety in $\overline{\Omega}$ on which each $f_{j}$ is holomorphic. We also compute the polynomial hull of the graph of pluriharmonic functions in some cases where the pluriharmonic functions are conjugates of holomorphic polynomials. We also give a couple of general theorem about uniform approximation on the domains with low boundary regularity.
Key words and phrases:
polynomial approximation; polynomial convexity, polynomial hull, pluriharmonic functions
2020 Mathematics Subject Classification:
Primary: 32E30, 32E20 Secondary: 46J10, 32A65
1. Introduction and statements of the results
Let $\mathcal{C}(K)$ represent the collection of all continuous complex-valued functions defined on a compact set $K$ in $\mathbb{C}^{n}$ with the norm $\|f\|_{K}=\sup_{x\in K}|f(x)|.$ For given complex-valued continuous functions $f_{1},\ldots,f_{N}$ on $K$, we use the notation $[z_{1},z_{2},\cdots,z_{n},f_{1},\ldots,f_{N};K]$ to represent the norm-closed subalgebra of $\mathcal{C}(K)$ that they generate. In $\mathbb{C}$, Wermer’s maximality theorem gives a very interesting dichotomy: A uniform algebra on the unit circle generated by holomorphic polynomials and some continuous functions $f_{1},\dots,f_{N}$ on $\mathbb{T}$ is either same as the uniform algebra generated by holomorphic polynomials restricted to the unit circle or the class of all continuous functions on the unit circle. That says either the algebra is of all continuous function or each of the function $f_{j}$ is holomorphic. Note that, by solving Dirichlet problem, any continuous function on the unit circle is boundary value of a harmonic function on the unit disc. In higher dimension, any continuous function on the Shilov boundary does not always extends to pluriharmonic functions. Pluriharmonic functions have holomorphic conjugates. Therefore, it is natural to consider functions that are boundary values of pluriharmonic functions to extend Wermer’s maximality theorem to higher dimensions. The study of uniform algebras generated by holomorphic and pluriharmonic functions in higher dimensions has been conducted by Čirka [32], Izzo [15, 16], Samuelsson and Wold [28], and Izzo, Samuelsson, and Wold [17]. In a seminal paper [28], Samuelsson and Wold proved the following result which extends Wermer’s maximality theorem for the case of bidisc in $\mathbb{C}^{2}$.
Result 1.1 (Samuelsson-Wold).
Let $f_{j}\in\mathcal{C}(\mathbb{T}^{2})$ for $j=1,\cdots,N,N\geq 1$ and assume that each $f_{j}$ extends to a pluriharmonic function on $\mathbb{D}^{2}.$ Then either $[z_{1},z_{2},f_{1},\cdots,f_{N};\mathbb{T}^{2}]=\mathcal{C}(\mathbb{T}^{2}),$ or there exists a non-trivial algebraic variety $Z\subset\mathbb{C}^{2}$ with $Z\cap b\mathbb{D}^{2}\subset\mathbb{T}^{2},$ and the pluriharmonic extensions of the $f_{j}$’s are holomorphic on $Z.$
Result 1.1 uses the properties of the bidisc crucially. One can ask the question: Does the same dichotomy hold when one considers other domains? This does not have immediate answer. For some images of the bidisc under proper polynomial maps, symmetrized bidisc being one of them, the question is answered positively in [11]. In this paper, we prove a similar result for certain polynomial polyhedra. Let $p_{1},p_{2},\cdots,p_{l}$ be holomorphic polynomials in $z_{1},z_{2},\ldots,z_{n}$. Consider the set $\mathfrak{D}_{l}$ known as the polynomial polyhedron defined by
$\mathfrak{D}_{l}:=\{z\in\mathbb{C}^{n}:|p_{1}(z)|<1,\ldots,|p_{l}(z)|<1\},$
where $l\geq n$. Define $\Sigma_{j}:=\{z\in\mathbb{C}^{n}:|p_{j}(z)|=1\}$ and $\Sigma_{J_{k}}:=\Sigma_{j_{1}}\cap\ldots\cap\Sigma_{j_{k}}$, where $J_{k}=\{j_{1},\ldots,j_{k}\}$ with $1\leq j_{1}<j_{2}<\ldots<j_{k}\leq l$. Clearly, the topological boundary $b\mathfrak{D}_{l}$ of $\mathfrak{D}_{l}$ is contained in $\bigcup_{j}\Sigma_{j}$, and following the definition from [6], we say $\mathfrak{D}_{l}$ is complex non-degenerate if
$dp_{j_{1}}\wedge\ldots\wedge dp_{j_{k}}(z)\neq 0\text{ on }\Sigma_{J_{k}},\quad 1\leq k\leq n.$
If $\mathfrak{D}_{l}$ is a complex non-degenerate analytic polyhedron, then the Shilov boundary coincides with the distinguished boundary $\Gamma_{\mathfrak{D}_{l}}$ of $\mathfrak{D}_{l}$ and is equal to $\Gamma_{\mathfrak{D}_{l}}=\bigcup^{\prime}\Sigma_{J_{n}}$, where $\bigcup^{\prime}$ represents the union over all monotone multi-indices $(j_{1},\cdots,j_{n})$. In particular, when the number of polynomials is $n$, then $\Gamma_{\mathfrak{D}_{n}}=\{z\in\mathbb{C}^{n}:|p_{j}(z)|=1,~{}j=1,\ldots,n\}$ (see [6], [12]). For function algebra on analytic polyhedron and Shilov boundary for uniform algebras on analytic polyhedron see [6] (see also [12, 13]), and for details on uniform algebras on arbitrary domain and Shilov boundary see [10].
Theorem 1.2.
Let $\mathfrak{D}_{2}$ be complex non-degenerate simply connected polynomial polyhedron given by $\overline{\mathfrak{D}}_{2}:=\{z\in\mathbb{C}^{2}:|p_{1}(z)|\leq 1,|p_{2}(z)|\leq 1\}.$ Assume that each leaf $\{z\in\mathbb{C}^{2}:\;p_{1}(z)=c\},$$\{z\in\mathbb{C}^{2}:\;p_{2}(z)=c\}$ is simply connected for $|c|=1.$ Let $f_{j}\in\mathcal{C}(\Gamma_{\mathfrak{D}_{2}})$ for $j=1,\cdots,N,N\geq 1,$ and assume that each $f_{j}$ extends to a pluriharmonic function on $\mathfrak{D}_{2}.$ Then either
$[z_{1},z_{2},f_{1},\cdots,f_{N};\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}})$
or there exists a non-trivial algebraic variety $V\subset\mathbb{C}^{2}$ with $V\cap b\mathfrak{D}_{2}\subset\Gamma_{\mathfrak{D}_{2}},$ and the pluriharmonic extension of each of the functions $f_{j}$ is holomorphic on $V.$
Remark 1.3.
If we take $p_{1}(z)=z_{1}$ and $p_{2}(z)=z_{2}$ we get back Samuelsson-Wold theorem (Result 1.1).
Remark 1.4.
Since each leaf is a simply connected Riemann surface, therefore, the number of components of each of the sets $\{z\in\mathbb{C}^{2}:\;p_{1}(z)=c\}$ and $\{z\in\mathbb{C}:\;p_{2}(z)=c\}$ is one for all $|c|=1.$
We also provide a description of the polynomial hull of certain graphs over the distinguished boundary of some polynomial polyhedra. In general, finding the polynomial hull is extremely difficult. Here we consider a polynomial polyhedron $\mathfrak{D}_{2}=\{z\in\mathbb{C}^{2}:\;|p_{1}(z)|\leq 1,|p_{2}(z)|\leq 1\}$ such that $\Psi:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by
$\Psi(z):=(p_{1}(z),p_{2}(z))$
is a proper map. Let $P$ be a holomorphic polynomial. We will compute the polynomially convex hull of the graph of $\overline{P\circ\Psi}$ over the distinguished boundary of the polynomial polyhedron $\mathfrak{D}_{2}$ (see Section 4 for details). Jimbo [18, 19] gave the description of polynomial hulls of graphs of anti-holomorphic polynomials on the torus. Here, we treat the case when $f$ is the restriction of an anti-holomorphic polynomial in $\mathbb{C}^{2}$ to the distinguished boundary $\Gamma_{{\mathfrak{D}_{2}}}$ of a complex non-degenerate polynomial polyhedron $\mathfrak{D}_{2}$ in $\mathbb{C}^{2}.$ We present explicit descriptions of the hull of the graph of $f$ on $\Gamma_{{\mathfrak{D}_{2}}}.$ In this setting, we show that the set $\widehat{{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})}\setminus{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})$ is contained in an algebraic variety. We now look at the graph over certain bounded domains whose closure is polynomially convex. Specifically, we look at domains with polynomially convex closure and having segment property.
Definition 1.5.
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}.$ $\Omega$ has segment property if for every $\xi\in b\Omega,$ there exists a neighborhood $N_{\xi}$ of $\xi$ and a vector $V_{\xi}$ in $\mathbb{C}^{n}$ such that
$z+tV_{\xi}\in\Omega,\forall z\in\overline{\Omega}\cap N_{\xi},0<t<1.$
In [4] authors showed that a domain $\Omega$ has the segment property if, and only if, $\Omega$ has $\mathcal{C}^{0}$ boundary. For more details on the segment property we refer the reader to [7]. We now present a couple of observations that will generalize results from ([28, Theorem 1.3]) and [17, Theorem 1.1] respectively:
Theorem 1.6.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain with $\widehat{\overline{{\Omega}}}=\overline{{\Omega}}$ and $\Omega$ possesses segment property except possibly at a countable set of boundary points. Let $h_{j}$’s be pluriharmonic on $\Omega$ and continuous on $\overline{\Omega}$ for $j=1,\cdots,N,$ and assume that $[z_{1},\cdots,z_{n},h_{1},\cdots,h_{N};\partial\Omega]=\mathcal{C}(\partial\Omega).$ Then either there exists an analytic disc in $\Omega$ where all $h_{j}$’s are holomorphic or $[z_{1},\cdots,z_{n},h_{1},\cdots,h_{N};\overline{\Omega}]=\mathcal{C}(\overline{\Omega}).$
Theorem 1.7.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain with $\widehat{\overline{{\Omega}}}=\overline{{\Omega}}$ and $\Omega$ possesses segment property except possibly at a countable set of boundary points. Let $h_{j}$’s be pluriharmonic on $\Omega$ and continuous on $\overline{\Omega}$ for $j=1,\cdots,N.$ Then the following are equivalent:
  1. (1)
    $G_{h}(\partial\Omega)$ is polynomially convex.
  2. (2)
    $\widehat{G_{h}(\partial\Omega)}\setminus G_{h}(\partial\Omega)$ contains no analytic disc.
  3. (3)
    There does not exist a nontrivial analytic disc $\triangle\to\Omega$ on which all the $h_{j}$’s are holomorphic.
  4. (4)
    $[z,h;\overline{\Omega}]=\left\{f\in\mathcal{C}(\overline{\Omega}):f|_{\partial\Omega}\in[z,h;\partial\Omega]\right\}.$
We observe that the $\mathcal{C}^{1}$-smoothness of $\Omega$ is required to prove ([28, Theorem 1.3]) and ([17, Theorem 1.1]) only to show that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex. By Lemma 2.21, we can say that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex for any bounded domain $\Omega$ with polynomially convex closure and having segment property except possibly at a countable set of boundary points. The rest follows from [28, Theorem 1.3] and [17, Theorem 1.1], respectively. We now give a brief outline for our proofs of Theorem 1.2. The philosophy behind the proof of Theorem 1.2 closely follows that of Result 1.1 although adapting the relevant tools to our case requires some new insights. For example, in the proof of Theorem 1.2, the polynomial convexity of $h$ over the polynomial polyhedron is crucial, which we prove by showing that the polynomial polyhedron has segment property. We also gave a version of Tornehave’s result for polynomial polyhedra. Firstly, we establish that Theorem 1.2 holds if and only if the graph of $h$ over the distinguished boundary of the polynomial polyhedron $\mathfrak{D}_{2}$ is polynomially convex (refer to Theorem 2.24). Therefore, if the graph lacks polynomial convexity, we examine different scenarios through specific lemmas, demonstrating the existence of an analytic variety within $\mathfrak{D}_{2}$ where all $h_{j}$’s are holomorphic. Utilizing Tornehave’s result (for polynomial polyhedron), we can find the required algebraic variety. A few words about the layout of the paper: The first part of Section 2 collects some known results that are required in this paper while in the second part we prove several new results which are crucial for this paper. In Section 3 we present a proof of Theorem 1.2. The polynomial hulls of some graphs are computed in Section 4. Section 5 is devoted to some examples.
2. Technical Results
We make use of the following result due to Samuelsson-Wold [28, Corollary 4.2].
Result 2.1 (Samuelsson-Wold).
Let $M$ be a complex manifold of dimension $n,$ let $h=(h_{1},\cdots,h_{N}):M\to\mathbb{C}^{N}$ be a $\mathcal{C}^{1}$-smooth map. If $z\in M$ and there are functions $h_{j_{1}},\cdots,h_{j_{n}}$ such that ¯$\overline{\partial}h_{j_{1}}\wedge\cdots\wedge\overline{\partial}h_{j_{n}}(z)\not=0,$ then the graph ${\sf Gr}_{h}(M)$ of $h$ at $(z,h(z))$ is totally real.
For $n>1,$ the most general result known for polynomial approximation is the following due to Samuelssonand Wold [28]:
Result 2.2.
(Samuelsson-Wold) Let $X$ be a polynomially convex compact set in $\mathbb{C}^{n}$ and assume that there are closed sets $X_{0}\subset X_{1}\subset\cdots\subset X_{N}=X$ such that $X_{j}\setminus X_{j-1},j=1,\cdots,N,$ is a totally real set. If $\mathcal{C}(X_{0})=\mathcal{O}(X_{0}),$ then $\mathcal{C}(X)=[z_{1},\cdots,z_{n};X].$
The following result shows that plurisubharmonic functions can also be used to describe the polynomial convex hull of a compact set.
Result 2.3 (Hörmander, [14]).
Let $G$ be a pseudoconvex open subset of $\mathbb{C}^{n}$ and $X$ be a compact subset $G.$ Then $\widehat{X}_{G}=\widehat{X}_{G}^{P}$, where $\widehat{X}_{G}^{P}:=\left\{z\in G:u(z)\leq\sup\nolimits_{X}u\;\forall u\in{\sf psh}(G)\right\}$ and $\widehat{X}_{G}:=\big{\{}z\in G:|f(z)|\leq\sup\nolimits_{z\in X}|f(z)|\;\forall f\in\mathcal{O}(G)\big{\}}$.
In case $G=\mathbb{C}^{n},$ Result 2.3 says that the plurisubharmonically convex hull of $X$ is same as the polynomially convex hull of $X.$
Result 2.4 ([25, 26]).
Let $E$ be a compact set in $\mathbb{C}^{n}.$ Let $a\in\widehat{E}\setminus E$, and let $N$ be a relatively compact neighborhood of $a$ that does not intersect $E$. Then, $a\in\widehat{{\widehat{E}\cap bN}}$ (where $bN$ denotes the topological boundary of $N$).
Result 2.5.
([28, Propostion 4.7]) Let $K\subset\mathbb{C}^{N}$ be compact, let $F:\mathbb{C}^{n}\to\mathbb{C}^{m}$ be the uniform limit on $K$ of entire functions, and let $Y=F(K);$ note that $F$ extends to $\widehat{K}.$ If $y\in Y$ is a peak point for the algebra $\mathscr{P}(Y),$ then $\widehat{K}\cap F^{-1}(y)=\widehat{(K\cap F^{-1}\{y\}}).$
Let ${\sf psh}(\Omega)$ denote the class of plurisubharmonic functions in $\Omega$ and ${\sf psh}(\overline{\Omega})$ denotes the collection of function $\phi$ such that there exists an open set $U\ \supset\Omega$ with $\phi\in{\sf psh}(U)$. Sibony [29, Theorem 2.2] showed that if $\Omega$ is a smooth pseudoconvex domain, then every functions in ${\sf psh}(\Omega)\cap C(\overline{\Omega})$ can be approximated uniformly with functions in ${\sf psh}(\overline{\Omega})\cap C^{\infty}(\overline{\Omega})$ and we say $\Omega$ has the PSH-Mergelyan property. Fornæss and Wiegerinck [9, Theorem 1] proved that this is also true for any bounded $\mathcal{C}^{1}$-smooth domain. Avelin, Hed, and Persson [5] further relaxed the boundary conditions to continuity. Subsequently, in [21], the authors presented an improved version of this result.
Result 2.6.
Let $\Omega$ be a bounded domain in $\mathbb{C}^{N}.$ Suppose that $\Omega$ has segment property except at a countable set of boundary points. Then $\Omega$ has the PSH-Mergelyan property.
The following definition is from [20].
Definition 2.7.
Two embeddings $\Phi,\Psi:\mathbb{C}\hookrightarrow\mathbb{C}^{n}$ are equivalent if there exist automorphisms $\phi\in\text{Aut}(\mathbb{C}^{n})$ and $\psi\in\text{Aut}(\mathbb{C})$ such that $\phi\circ\Phi=\Psi\circ\psi.$
Result 2.8.
Every polynomial embedding of $\mathbb{C}$ into $\mathbb{C}^{2}$ is equivalent to the standard one, i.e. if a complex algebraic curve in $\mathbb{C}^{2}$ is isomorphic to $\mathbb{C}$ then in some algebraic coordinates $(z^{*}_{1},z^{*}_{2})$ on $\mathbb{C}^{2}$ its equation is $z^{*}_{1}=0$ (see Abhyankar-Moh [1], and Suzuki [31]).
The following result [27, Theorem 5.1] gives a criterion for a homogeneous polynomial map from $\mathbb{C}^{n}$ to $\mathbb{C}^{n}$ to be proper holomorphic.
Result 2.9 (Rudin).
Suppose that
  1. (1)
    $d_{1},\cdots,d_{n}$ are positive integers,
  2. (2)
    $p_{j}:\mathbb{C}^{n}\to\mathbb{C}$ is a homogeneous polynomial of degree $d_{j},$ for $j=1,\cdots,n,$
  3. (3)
    Define $F:\mathbb{C}^{n}\to\mathbb{C}^{n}$ by $F(z_{1},\cdots,z_{n})=(p_{1}(z_{1},\cdots,z_{n}),\cdots,p_{n}(z_{1},\cdots,z_{n})).$
Then $F:\mathbb{C}^{n}\to\mathbb{C}^{n}$ is proper if and only if $F^{-1}\{0\}=\{0\}.$
The following result is due to Izzo-Samuelsson-Wold [17, Lemma 2.2].
Result 2.10 (Izzo-Samuelsson-Wold).
Let $\Omega\subset\mathbb{C}^{n}$ be a domain, let $h_{j}\in PH(\Omega)$ for $j=1,\cdots,N,$ and $Z\subset\Omega$ be an irreducible analytic set of dimension $d\geq 1.$ Let $1\leq m\leq d,$ fix $(i_{1},\cdots,i_{m}),$ and define
[…]
Then $\widetilde{Z}:=Z^{{}^{\prime}}\cup Z_{sing}$ is an analytic subset of $\Omega.$
The following result is from [2, Lemma 1.3].
Result 2.11 (Alexander).
If $X$ is a compact subset of $\mathbb{C}^{n}$ and $E$ is a component of $\widehat{X}$, then $E=\widehat{X\cap E}$ (by a component of $X,$ we mean a maximal connected subset of $X$).
The following result is known as the Remmert proper mapping theorem. It will play a crucial role in our proofs.
Result 2.12 (Remmert Proper Mapping theorem [23, 24]).
Let $M,N$ be complex spaces and $f:M\to N$ is a proper holomorphic map. If $Z$ is an analytic subvariety in $M$ then $f(Z)$ is also an analytic subvariety in $N.$ Moreover, if $Z$ is irreducible then $f(Z)$ is also irreducible subvariety of $N.$
The following is known as Tornehave’s result ([30, Corollary 3.8.11]).
Result 2.13 (Tornehave).
Let $V$ be a one-dimensional analytic subvariety of $\mathbb{D}^{n}$ such that $\overline{V}\setminus V\subset\mathbb{T}^{n}.$ Then there is an algebraic curve $Z$ in $\mathbb{C}^{n}$ with $V=Z\cap\mathbb{D}^{n}.$
The following is a very simple lemma.
Lemma 2.14.
Let $h$ be a complex valued pluriharmonic function on a simply connected domain $\Omega,$ then there exist holomorphic functions $\Psi$ and $\Phi$ on $\Omega$ such that $h=\Psi+{\sf Re}\Phi$ on $\Omega.$
In the remaining part of this section we state and proof some results which are mostly new. These will be used in the proofs of our theorems.
Lemma 2.15.
Let $\mathfrak{D}_{2}=\{z\in\mathbb{C}^{2}:|p_{1}(z)|<1,|p_{2}(z)|<1\}$ be a bounded complex non-degenerate polynomial polyhedron and $f\in PH(\mathfrak{D}_{2})\cap\mathcal{C}(\overline{\mathfrak{D}_{2}}).$ Then $f$ is harmonic on $b\mathfrak{D}_{2}.$
Proof.
Let $\alpha\in\mathcal{A}:=\{|p_{1}(z)|=1,|p_{2}(z)|<1\}\subset b\mathfrak{D}_{2}.$ Since $\mathfrak{D}_{2}$ is complex non-degenerate, $\mathfrak{D}_{2}$ has segment property. Therefore, there exists a vector $V_{\alpha}$ and a neighborhood $N(\alpha)$ of $\alpha$ in $\mathbb{C}^{2}$ such that
$z+tV_{\alpha}\in\mathfrak{D}_{2},\forall z\in N(\alpha)\cap\overline{\mathfrak{D}_{2}},\forall t\in(0,1).$
We choose an analytic disc $\phi:\mathbb{D}\to\mathbb{C}^{2}$ in $N(\alpha)\cap\mathcal{A}$ i.e., $\phi(\mathbb{D})\subset N(\alpha)\cap\mathcal{A}.$ Since harmonic is a local property, it is enough to show that $f\circ\phi:\mathbb{D}\to\mathbb{C}$ is harmonic. We define a sequence of analytic disc $\phi_{n}$ by
$\phi_{n}(\xi)=\phi(\xi)+\frac{1}{n}V_{\alpha},\forall n\in\mathbb{N}.$
Since $\phi(\xi)\in(N(\alpha)\cap\mathcal{A})\subset N(\alpha)\cap\overline{\mathfrak{D}_{2}},\forall\xi\in\mathbb{D},$ hence $\phi(\xi)+\frac{1}{n}V_{\alpha}\in\mathfrak{D}_{2}$ i.e., $\phi_{n}(\mathbb{D})\subset\mathfrak{D}_{2}~{}\forall n\in\mathbb{N}.$ Since $f\in PH(\mathfrak{D}_{2})\cap\mathcal{C}(\overline{\mathfrak{D}_{2}}),$ $f\circ\phi_{n}:\mathbb{D}\to\mathbb{C}$ is harmonic for all $n.$ It is easy to see that $\phi_{n}$ converges to $\phi$ uniformly on $\mathbb{D}$. Indeed,
$\lim_{n\to\infty}\sup_{\mathbb{D}}\|\phi_{n}(\xi)-\phi(\xi)\|=\lim_{n\to\infty}\frac{1}{n}\|V_{\alpha}\|=0.$
Therefore, $f\circ\phi_{n}$ is converges to $f\circ\phi$ uniformly on $\mathbb{D}.$ Being uniform limit of sequence of harmonic functions, $f\circ\phi$ is also harmonic. ∎
Remark 2.16.
Let $\mathfrak{D}_{n}=\{z\in\mathbb{C}^{n}:|p_{1}(z)|<1,\cdots,|p_{n}(z)|<1\}$ be a bounded complex non-degenerate polynomial polyhedron and $f\in PH(\mathfrak{D}_{n})\cap\mathcal{C}(\overline{\mathfrak{D}_{n}}).$ Then for every increasing multi-index $J_{k}$ of length $k\in\{1,\cdots,n-1\}$ the function $f$ is pluriharmonic on the $(n-k)-$dimensional complex manifold $\Sigma_{J_{k}}.$
Next, we give a class of domains that has the segment property. Let $\psi_{1},\cdots,\psi_{N}:\mathbb{C}^{n}\to\mathbb{R}$ are $C^{1}$-smooth functions. Consider the bounded domain
$\Omega:=\{z\in\mathbb{C}^{n}:\psi_{1}<0,\cdots,\psi_{N}<0\}.$
We denote $\Sigma_{j}:=\{z\in\mathbb{C}^{n}:\psi_{j}(z)=0\}$ and $\Sigma_{J_{k}}:=\Sigma_{j_{1}}\cap\cdots\cap\Sigma_{j_{k}}$, where $J_{k}=({j_{1},\cdots,j_{k}})$ with $1\leq j_{1}<j_{2}<\ldots<j_{k}\leq N$. The topological boundary of $\Omega$ is denoted by $\partial\Omega$ and clearly $\partial\Omega\subset\cup_{j}\Sigma_{j}.$ The result presented below describes a class of domains where the segment property holds. A special case for Weil polyhedra was proved in [22, Proposition 2.1].
Proposition 2.17.
Let $\Omega$ be as above. Assume that for any collection $\{j_{1},\cdots j_{k}\}\subset\{1,2,\cdots,N\},$ $d\psi_{j_{1}}\wedge\cdots\wedge d\psi_{j_{k}}(z)\neq 0\text{ on }\Sigma_{J_{k}}~{}~{}\forall~{}1\leq k\leq N.$ Then $\Omega$ has the segment property.
Proof.
Let $\xi\in\partial\Omega.$ This implies $\xi\in\Sigma_{J_{k}}$ for some $J_{k}$ i.e., there exist $j_{1},\cdots,j_{k}$ such that
$\psi_{j}(\xi)$ $=0~{}~{}\forall j\in\{j_{1},\cdots,j_{k}\}\text{ and }$
$\psi_{s}(\xi)$ $<0~{}~{}\forall s\notin\{j_{1},\cdots,j_{k}\}.$
Let us define $\Psi_{J_{k}}:=(\psi_{j_{1}},\cdots,\psi_{j_{k}}):\mathbb{C}^{n}\to\mathbb{R}^{k},$ where $J_{k}=({j_{1},\cdots,j_{k}}).$ Then the differential $d\Psi_{J_{k}}$ of $\Psi_{J_{k}}$ at $\xi$ is given by
$d\Psi_{J_{k}}|_{\xi}=\begin{pmatrix}\frac{\partial\psi_{j_{l}}(\xi)}{\partial{z}_{m}}\vline\frac{\partial\psi_{j_{l}}(\xi)}{\partial{\bar{z}}_{m}}\end{pmatrix}^{m=1,\cdots,n}_{l=1,\cdots,k}$
The condition $d\psi_{j_{1}}\wedge\cdots\wedge d\psi_{j_{k}}(z)\neq 0\text{ on }\Sigma_{J_{k}}$ implies that $d\Psi_{J_{k}}$ has maximal rank on $\Sigma_{J_{k}}.$ Let $\widetilde{V}=(\widetilde{V_{1}},\cdots,\widetilde{V_{k}})\in\mathbb{R}^{k}$ such that $\widetilde{V_{j}}<0~{}~{}\forall j\in\{1,\cdots,k\}.$ Then there exists $W=(W_{1},\cdots,W_{n})\in\mathbb{C}^{n}$ such that
$d\Psi_{J_{k}}|_{\xi}(W)=\begin{pmatrix}\frac{\partial\psi_{j_{l}}(\xi)}{\partial{z}_{m}}\vline\frac{\partial\psi_{j_{l}}(\xi)}{\partial{\bar{z}}_{m}}\end{pmatrix}\begin{pmatrix}W\\ \overline{W}\end{pmatrix}=\widetilde{V}.$ (2.4)
If we take $w=W+\xi,$ where $w=(w_{1},\cdots,w_{n}),$ then from (2.4), we get that
$\sum_{l=1}^{n}\frac{\partial\psi_{j}(\xi)}{\partial{{z_{l}}}}({w_{l}}-{\xi_{l}})+\sum_{l=1}^{n}\frac{\partial\psi_{j}(\xi)}{\partial\bar{{z_{l}}}}(\bar{w_{l}}-\bar{\xi_{l}})$ $<0~{}~{}\forall j\in\{j_{1},\cdots,j_{k}\}.$
This implies
${\sf Re}\left<\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>$$<0,\text{ where }\partial\psi_{j}(\xi)=\left(\frac{\partial\psi_{j}(\xi)}{\partial{z}_{1}},\cdots,\frac{\partial\psi_{j}(\xi)}{\partial{z}_{n}}\right).$ (2.5)
Next we claim that there exists a neighborhood $U(\xi)$ of $\xi$ such that
${\sf Re}\left<\partial\psi_{j}(z),\bar{w}-\bar{\xi}\right>$$<0~{}~{}~{}~{}\forall z\in U(\xi)\text{ and }j\in\{j_{1},\cdots,j_{k}\}.$
Fix $j\in\{j_{1},\cdots,j_{k}\}.$ Using (2.5), for fixed $\xi$ and $w,$ we can choose $\varepsilon_{0,j}$ such that
${\sf Re}\left<\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>+\varepsilon_{0,j}<0.$ (2.6)
We write
${\sf Re}\left<\partial\psi_{j}(z),\bar{w}-\bar{\xi}\right>$ $={\sf Re}\left<\partial\psi_{j}(z)-\partial\psi_{j}(\xi)+\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>$
$={\sf Re}\left<\partial\psi_{j}(z)-\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>+{\sf Re}\left<\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>.$ (2.7)
Since $\partial\psi_{j}$ is continuous, for $\varepsilon=\frac{\varepsilon_{0,j}}{\|w-\xi\|}$ there exists $\delta_{1,j}>0$ such that
$\|\partial\psi_{j}(z)-\partial\psi_{j}(\xi)\|<\frac{\varepsilon_{0,j}}{\|w-\xi\|}~{}~{}~{}\forall z\text{ satisfying }0<\|z-\xi\|<\delta_{1,j}.$ (2.8)
Using Cauchy-Schwarz inequality together with (2.8), we get that
${\sf Re}\left<\partial\psi_{j}(z)-\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>$$\|\left<\partial\psi_{j}(z)-\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>\|<\frac{\varepsilon_{0,j}}{\|w-\xi\|}\|w-\xi\|<\varepsilon_{0,j}.$
This implies,
${\sf Re}\left<\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>+{\sf Re}\left<\partial\psi_{j}(z)-\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right>$$\varepsilon_{0,j}+{\sf Re}\left<\partial\psi_{j}(\xi),\bar{w}-\bar{\xi}\right><0.$
Hence
${\sf Re}\left<\partial\psi_{j}(z),\bar{w}-\bar{\xi}\right><0~{}\text{ for }\|z-\xi\|<\delta_{1,j}.$
Let us now define $U(\xi):=\bigcap_{j\in\{j_{1},\cdots,j_{k}\}}B_{\delta_{1,j}}(\xi),$ where $B_{\delta_{1,j}}(\xi):=\{z\in\mathbb{C}^{n}:\|z-\xi\|<\delta_{1,j}\}.$ Then for all $j\in\{j_{1},\cdots,j_{k}\},$ we obtain
$\sum_{l=1}^{n}\left[\frac{\partial\psi_{j}(z)}{\partial{{z_{l}}}}({w_{l}}-{\xi_{l}})+\frac{\partial\psi_{j}(z)}{\partial\bar{{z_{l}}}}(\bar{w_{l}}-\bar{\xi_{l}})\right]<0,~{}\forall z\in U(\xi).$ (2.9)
We take $W_{\xi}=(w-\xi).$ Now for $z\in\overline{U(\xi)}$ and $\eta>0,$ we get, by expanding in Taylor series
$\psi_{j}(z+\eta W_{\xi})=\psi_{j}(z)+\eta\left[\sum_{l=1}^{n}\frac{\partial\psi_{j}(z)}{\partial{{z_{l}}}}({w_{l}}-{\xi_{l}})+\frac{\partial\psi_{j}(z)}{\partial\bar{{z_{l}}}}(\bar{w_{l}}-\bar{\xi_{l}})\right]+g_{j}(\eta),$
where $g_{j}(\eta)=o(\eta).$ We define
$F_{j}(z):=\sum_{l=1}^{n}\frac{\partial\psi_{j}(z)}{\partial{{z_{l}}}}({w_{l}}-{\xi_{l}})+\frac{\partial\psi_{j}(z)}{\partial\bar{{z_{l}}}}(\bar{w_{l}}-\bar{\xi_{l}}).$
Therefore,
$\psi_{j}(z+\eta W_{\xi})=\psi_{j}(z)+\eta F_{j}(z)+g_{j}(\eta).$ (2.10)
Shrinking $U(\xi),$ if needed, we get
$\psi_{j}(z)$ $\leq 0~{}~{}\forall z\in\overline{\Omega},$
and by (2.9), we have
$F_{j}(z)$ $<0,~{}\forall z\in\overline{U(\xi)}.$
Since $\overline{U(\xi)}$ is compact, there exists $\varepsilon>0$ such that $F_{j}(z)+\varepsilon<0.$ For this $\varepsilon,$ we can find $\eta_{0,j}>0$ such that $|g_{j}(\eta)|<\varepsilon\eta_{0,j}.$ Therefore, from (2.10), we get that
$\psi_{j}(z+\eta W_{\xi})\leq\psi_{j}(z)+\eta_{0,j}[F_{j}(z)+\varepsilon]<0,~{}\forall\eta<\eta_{0,j}.$
We take $\tilde{\eta}_{1}:=\min\{\eta_{0,j_{1}},\cdots,\eta_{0,j_{k}}\}.$ Then we have
$\psi_{j}(z+\eta W_{\xi})<0,~{}\forall\eta<\tilde{\eta}_{1},\forall j\in\{j_{1},\cdots,j_{k}\},\text{ and }\forall z\in\overline{U(\xi)}\cap\overline{\Omega}.$
Next, let us fix $s\notin\{j_{1},\cdots,j_{k}\}.$ Then $\psi_{s}(\xi)<0.$ By continuity of $\psi_{s},$ there exists a neighborhood $B_{s}(\xi)$ of $\xi$ such that
$\psi_{s}(z)<0,~{}\forall z\in B_{s}(\xi).$
For $z\in B_{s}(\xi)$ and for some $\delta_{1}>0,$ we have
$\psi_{s}(z+\delta_{1}W_{\xi})=\psi_{s}(z)+$
$\delta_{1}\left[\sum_{l=1}^{n}\frac{\partial\psi_{s}(z)}{\partial{{z_{l}}}}({w_{l}}-{\xi_{l}})+\frac{\partial\psi_{s}(z)}{\partial\bar{{z_{l}}}}(\bar{w_{l}}-\bar{\xi_{l}})\right]+q_{s}(\delta_{1}),\text{ where }q_{s}(\delta_{1})=o(\delta_{1}).$ (2.11)
Let us define
$G_{s}(z):=\left[\sum_{l=1}^{n}\frac{\partial\psi_{s}(z)}{\partial{{z_{l}}}}({w_{l}}-{\xi_{l}})+\frac{\partial\psi_{s}(z)}{\partial\bar{{z_{l}}}}(\bar{w_{l}}-\bar{\xi_{l}})\right].$
Therefore,
$\psi_{s}(z+\delta_{1}W_{\xi})=\psi_{s}(z)+\delta_{1}G_{s}(z)+q_{s}(\delta_{1}).$
Since $G_{s}$ is continuous, there exists an $M>0$ such that
$G_{s}(z)\leq M~{}~{}~{}\forall z\in\overline{B_{s}(\xi)}.$
Therefore, (2) becomes
$\psi_{s}(z+\delta_{1}W_{\xi})\leq\psi_{s}(z)+\delta_{1}M+q_{s}(\delta_{1}).$
Note that $\psi_{s}<0$ on $\overline{B_{s}(\xi)}$ and $\delta_{1}M+q_{s}(\delta_{1})\to 0$ as $\delta_{1}\to 0.$ Choose $\varepsilon>0$ such that $\psi_{s}(z)+\varepsilon<0$ on $\overline{B_{s}(\xi)}.$ For this $\varepsilon>0,$ there exists $\delta_{0,s}>0$ such that
$|\delta_{1}M+q_{s}(\delta_{1})|<\varepsilon~{}~{}~{}~{}\text{ for}\delta_{1}<\delta_{0,s}.$
Take $B(\xi):=\displaystyle\bigcap_{s\notin\{j_{1},\cdots,j_{k}\}}B_{s}(\xi).$ Therefore, for all $z\in\overline{B(\xi)},$ $\psi_{s}(z+\delta_{1}W_{\xi})<0$ for $\forall\delta_{1}<\delta_{0,s}.$ Choose $\tilde{\eta}_{2}:=\min\{\delta_{0,s}:s\notin\{j_{1},\cdots,j_{k}\}\},$ and we define $\gamma:=\min\{\tilde{\eta}_{1},\tilde{\eta}_{2}\}.$ Therefore,
[…]
If we take $V_{\xi}:=\gamma W_{\xi},$ then we get from above that
$\psi_{j}(z+tV_{\xi})<0~{}~{}~{}~{}~{}~{}\forall t,0<t<1~{}~{}~{}~{}\text{ and }\forall z\in N({\xi})\cap\overline{\Omega}.$
Hence, $\Omega$ has the segment property. ∎
Corollary 2.18.
Let $\mathfrak{D}_{N}$ be a bounded complex non-degenerate polynomial polyhedron. Then $\mathfrak{D}_{N}$ has the segment property.
Proof.
We define $\psi_{j}(z)=|p_{j}(z)|^{2}-1:\mathbb{C}^{n}\to\mathbb{R}~{}~{}\forall j=1,\cdots,N,N\geq n.$ Then
$\mathfrak{D}_{N}:=\{z\in\mathbb{C}^{n}:\psi_{1}<0,\cdots,\psi_{N}<0\}.$
Since $\mathfrak{D}_{N}$ is a complex non-degenerate polynomial polyhedron, $dp_{j_{1}}\wedge\cdots\wedge dp_{j_{k}}(z)\neq 0\text{ on }\Sigma_{J_{k}}~{}~{}\forall~{}1\leq k\leq N.$ Therefore, $d\psi_{j_{1}}\wedge\cdots\wedge d\psi_{j_{k}}(z)\neq 0\text{ on }\Sigma_{J_{k}}~{}~{}\forall~{}1\leq k\leq N.$ Now by using Proposition 2.17, we can say that $\mathfrak{D}_{N}$ has the segment property. ∎
Lemma 2.19.
Suppose that $d_{1},\cdots,d_{n}$ are positive integers and $p_{j}:\mathbb{C}^{n}\to\mathbb{C}$ is a homogeneous polynomial of degree $d_{j},$ for $j=1,\cdots,n.$ Then the polynomial polyhedron $\overline{\mathfrak{D}}_{n}$ given by $\mathfrak{D}_{n}:=\{z\in\mathbb{C}^{n}:|p_{1}(z)|<1,\cdots,|p_{n}(z)|<1\}$ is contractible. Additionally, if we take $F=(p_{1},\cdots,p_{n})$ and $F^{-1}\{0\}=\{0\},$ then $\overline{\mathfrak{D}}_{n}:=\{z\in\mathbb{C}^{n}:|p_{1}(z)|\leq 1,\cdots,|p_{n}(z)|\leq 1\}$ is compact.
Proof.
Since each $p_{j}$ is a homogeneous polynomial, $0=(0,\cdots,0)\in\mathfrak{D}_{n}.$ Let $t\in[0,1]$ and $z\in\mathfrak{D}_{n}.$ Then $|p_{j}(tz)|=|t^{d_{j}}p_{j}(z)|\leq|p_{j}(z)|<1.$ Therefore, $\mathfrak{D}_{n}$ is star-shaped with respect to the origin. We know that every star-shaped domain is contractible (via straight-line homotopy).
For the second part, assume $F=(p_{1},\cdots,p_{n})$ and $F^{-1}\{0\}=\{0\}.$ By Result 2.9, it follows that $F$ is a proper holomorphic map. Since $F^{-1}(\overline{\mathbb{D}}^{n})=\overline{\mathfrak{D}}_{n},$ it follows that $\overline{\mathfrak{D}}_{n}$ is compact, where $\overline{\mathbb{D}}^{n}$ is the closed unit polydisc in $\mathbb{C}^{n}.$
Note that every closed bounded connected polynomial polyhedron in $\mathbb{C}$ is simply connected. However, the following example illustrates that every closed bounded polynomial polyhedron in $\mathbb{C}^{2}$ may not be simply connected.
Example 2.20.
Let $\overline{\mathfrak{D}}_{2}:=\{z\in\mathbb{C}^{2}:|z_{1}z_{2}-1|\leq\frac{1}{2},|z_{1}+z_{2}-1|\leq 4\}.$ Then $\mathfrak{D}_{2}$ is not simply connected. Note that $\overline{\mathfrak{D}}_{2}\subset\{z\in\mathbb{C}^{2}:z_{1}\neq 0\}.$ Let $\gamma:=\{(e^{i\theta},e^{-i\theta}):0\leq\theta\leq 2\pi\}$ be a loop in $\overline{\mathfrak{D}}_{2}$ and take the projection map $\Pi:\{z\in\mathbb{C}^{2}:z_{1}\neq 0\}\to\mathbb{C}^{*}$ by $(z_{1},z_{2})\to z_{1}.$ Clearly, $\Pi(\gamma)=\{e^{i\theta}:0\leq\theta\leq 2\pi\}$ is not null homotopic in $\mathbb{C}^{*}.$ Hence, $\gamma$ is not null homotopic in $\overline{\mathfrak{D}}_{2}.$ Therefore, $\overline{\mathfrak{D}}_{2}$ is not simply connected.
The proof of the following result for a $C^{1}$-smooth domain is given in Samuelsson and Wold [28, Lemma 4.5].
Lemma 2.21.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain with $\overline{\Omega}$ polynomially convex. Also, assume that $\Omega$ has the segment property except possibly at a countable set of boundary points. Let $h_{j}\in PH(\Omega)\cap\mathcal{C}(\overline{\Omega})$ for $j=1,\cdots,N.$ Then ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex.
Proof.
Since $\overline{\Omega}$ is polynomially convex, $\widehat{{\sf Gr}_{h}(\overline{\Omega})}\subset\overline{\Omega}\times\mathbb{C}^{N}.$ Let $(\alpha,\beta)\in\left(\overline{\Omega}\times\mathbb{C}^{N}\right)\setminus{\sf Gr}_{h}(\overline{\Omega}).$ Therefore, $\beta\neq h(\alpha).$ Then there exists $l\in\{1,\cdots,N\}$ such that $h_{l}(\alpha)\neq\beta_{l}.$ We assume that ${\sf Re}(h_{l}(\alpha)-\beta_{l})>0$. Let us define $\Psi_{l}(z,w)={\sf Re}h_{l}(z)-{\sf Re}w_{l}.$ Clearly, ${\sf Re}h_{l}(z)$ is plurisubharmonic function on $\Omega$ and continuous on $\overline{\Omega}$. By Result 2.6, the function ${\sf Re}h_{l}$ can be approximated uniformly on $\overline{\Omega}$ by a sequence of functions $\{\eta_{j}\}$ in ${\sf psh}(\overline{\Omega})\cap\mathcal{C}^{\infty}(\overline{\Omega}).$ If $\tilde{\eta}_{j}:=\eta_{j}-{\sf Re}w_{l},$ then $\{\tilde{\eta}_{j}\}$ approximates $\Psi_{l}$ uniformly on compact subsets of $\overline{\Omega}\times\mathbb{C}^{N}.$ Moreover, we have $\Psi_{l}(\alpha,\beta)>\sup_{{\sf Gr}_{h}(\overline{\Omega})}\Psi_{l}=0.$ Since $\overline{\Omega}$ is polynomially convex, there is an open Runge and Stein neighborhood $\widetilde{\Omega}\supset\overline{\Omega}$ and a function $\widetilde{\Psi_{l}}\in{\sf psh}(\widetilde{\Omega}\times\mathbb{C}^{N})$ such that $\widetilde{\Psi_{l}}(\alpha,\beta)>\sup_{{\sf Gr}_{h}(\overline{\Omega})}\widetilde{\Psi_{l}}.$ Using Hörmander’s result (Result 2.3), we get that $(\alpha,\beta)\notin\widehat{{\sf Gr}_{h}(\overline{\Omega})}_{\mathcal{O}(\widetilde{\Omega}\times\mathbb{C}^{N})},$ and hence $(\alpha,\beta)\notin\widehat{{\sf Gr}_{h}(\overline{\Omega})}.$ Thus ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex. ∎
Next, we state and prove a version of $Tornehave$’s result for polynomial polyhedron.
Lemma 2.22.
Let the polynomials $p_{1},\cdots,p_{n}$ define the polynomial polyhedron $\mathfrak{D}_{n}$, and $\Psi=(p_{1},\cdots,p_{n})$ is a proper polynomial map. Let $V$ be an irreducible analytic variety in $\mathfrak{D}_{n}$ of dimension one such that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{n}}.$ Then there exists an algebraic variety $\mathcal{Z}\subset\mathbb{C}^{n}$ of dimension one such that $V=\mathcal{Z}\cap\mathfrak{D}_{n}.$
Proof.
Let $V$ be an irreducible analytic variety of dimension one in $\mathfrak{D}_{n}$ such that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{n}}.$ By Result 2.12, $\Psi(V)$ is also an irreducible analytic variety of dimension one in $\mathbb{D}^{n}.$ It is easy to show that $\overline{\Psi(V)}\setminus\Psi(V)\subset\mathbb{T}^{n}.$ Then by Result 2.13, there exists an algebraic variety $Z\subset\mathbb{C}^{n}$ such that $\Psi(V)=Z\cap\mathbb{D}^{n}.$ Therefore, $\Psi^{-1}(\Psi(V))=\Psi^{-1}(Z\cap\mathbb{D}^{n})=\Psi^{-1}(Z)\cap\Psi^{-1}(\mathbb{D}^{n}).$ Since $Z$ is an algebraic variety in $\mathbb{C}^{n}$ and $\Psi$ is a polynomial map, $\Psi^{-1}(Z)$ is also an algebraic variety in $\mathbb{C}^{n}.$ Assume that $\Psi^{-1}(Z)=\cup Z_{j},$ where $Z_{j}$ is an irreducible component of $\Psi^{-1}(Z)$ of dimension one. Clearly, $V\subset\Psi^{-1}(Z)=\cup Z_{j}.$ We claim that $V\subset Z_{j_{0}}$ for some $j_{0}.$ Without loss of generality, we assume that $V\nsubseteq Z_{1},V\nsubseteq Z_{2}$ but $V\subset Z_{1}\cup Z_{2}.$ Hence $V=V\cap(Z_{1}\cup Z_{2})=Z_{1}\cap V\cup Z_{2}\cap V.$ Therefore, $V\setminus(Z_{1}\cap Z_{2})=(Z_{1}\cap V\cup Z_{2}\cap V)\setminus(Z_{1}\cap Z_{2}).$ This implies,
$V\setminus(Z_{1}\cap Z_{2})=((Z_{1}\cap V)\setminus(Z_{1}\cap Z_{2}))\cup((Z_{2}\cap V)\setminus(Z_{1}\cap Z_{2})).$ (2.12)
We set, $S:=V_{Sing}\cap(Z_{1}\cap Z_{2}).$ Note that every point of $Z_{1}\cap Z_{2}$ is a singular point for $Z_{1}\cup Z_{2}$, and hence isolated and regular point of the variety $V$ can not be isolated. Here $V$ is one-dimensional irreducible variety. Therefore, no point of $Z_{1}\cap Z_{2}$ is a regular point of $V.$ Hence we obtain that, $V\setminus S=V\setminus(Z_{1}\cap Z_{2}).$ Therefore, from (2.12) we get that
$V\setminus S=((Z_{1}\cap V)\setminus(Z_{1}\cap Z_{2}))\cup((Z_{2}\cap V)\setminus(Z_{1}\cap Z_{2})).$ (2.13)
Since $S\subset V_{Sing}$ and the set of regular points of an irreducible variety is connected, therefore we can say that $V\setminus S$ is connected. But right hand side of (2.13) is disconnected, and this is a contradiction to the assumption that $V\nsubseteq Z_{1},V\nsubseteq Z_{2}$ but $V\subset Z_{1}\cup Z_{2}.$ Without loss of generality, we assume that $V\subset Z_{1}.$ Next, we claim that $V=\mathfrak{D}_{n}\cap Z_{1}.$ Assume that $V\subset Z_{1}\cap\mathfrak{D}_{n}.$ Since $Z_{1}$ is an irreducible variety and if $V$ is proper subset of $Z_{1},$ then $\dim_{\mathbb{C}}V<\dim_{\mathbb{C}}Z_{1}$ (see Izzo [3], Federer [8, Theorem 3.4.8, assertion (15)]). This is not possible, as otherwise, $V$ will be discrete. Therefore, $V=\mathfrak{D}_{n}\cap Z_{1}.$
The proof of the following result for a compact connected subset of $\mathbb{C}^{n}$ is given in Jimbo [18, Main Lemma].
Lemma 2.23.
Let $X$ be compact subset of $\mathbb{C}^{n}.$ Let $\Omega$ be an open subset of $\mathbb{C}^{n}$ such that $X\cap\Omega=\emptyset.$ If $\widehat{X}\cap\Omega$ is contained in a totally real manifold $M$ of $\Omega,$ then $\widehat{X}\cap\Omega=\emptyset.$
Theorem 2.24.
Let $p_{1},\cdots,p_{n}$ be holomorphic polynomials in $\mathbb{C}^{n}$ such that $z\longmapsto(p_{1}(z),\cdots,p_{n}(z))$ is a proper holomorphic map from $\mathbb{C}^{n}$ to $\mathbb{C}^{n}.$ Define $\Gamma:=\{z\in\mathbb{C}^{n}:|p_{1}(z)|=1,\cdots,|p_{n}(z)|=1\}$ and assume that $dp_{1}\wedge\cdots\wedge dp_{n}(z)\neq 0\text{ on }\Gamma.$ Let $N\geq 1$ and $f_{1},\cdots,f_{N}\in\mathcal{C}(\Gamma).$ Then $[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]=\mathcal{C}(\Gamma)$ if and only if ${\sf Gr}_{f}(\Gamma)$ is polynomially convex, where $f=(f_{1},\cdots,f_{N}).$
Proof.
Since the map $z\longmapsto(p_{1}(z),\cdots,p_{n}(z))$ is proper map from $\mathbb{C}^{n}$ to $\mathbb{C}^{n},$ $\Gamma$ is compact. First, we show that $[z_{1},\cdots,z_{n},\overline{{p_{1}}},\cdots,\overline{{p_{n}}};\Gamma]=\mathcal{C}(\Gamma).$ We consider
$K:$ =$=\{(z_{1},\cdots,z_{n},w_{1},\cdots,w_{n}):w_{j}=\overline{p_{j}(z)},|p_{j}(z)|=1,j=1,\cdots,n\}$
=$=\{(z_{1},\cdots,z_{n},w_{1},\cdots,w_{n}):w_{j}p_{j}(z)=1,|p_{j}(z)|=1,j=1,\cdots,n\}\subset\mathbb{C}^{2n}.$
We claim that $K$ is polynomially convex. To prove this claim, we assume that $(a,b)\in\mathbb{C}^{2n}\setminus K,$ where $a=(a_{1},\cdots,a_{n})$ and $b=(b_{1},\cdots,b_{n}).$ Case I: $b_{l}p_{l}(a)\not=1,$ $|p_{l}(a)|=1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=w_{l}p_{l}(z)-1.$ Then
$|\psi(a,b)|>0=\sup_{K}|\psi|.$
Hence, in this case, $(a,b)\notin\widehat{K}.$ Case II: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|<1$ for some $l\in\{1,\cdots,n\}.$ This implies $|b_{l}|>1.$ Consider the polynomial $\psi(z,w)=w_{l}.$ Then
$|b_{l}|=|\psi(a,b)|>1=\sup_{K}|\psi|.$
Hence, in this case, $(a,b)\notin\widehat{K}.$ Case III: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|>1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=p_{l}(z).$ Then
$|p_{l}(a)|=|\psi(a,b)|>1=\sup_{K}|\psi|.$
Hence, in this case, $(a,b)\notin\widehat{K}.$ Combining all three cases, we conclude that $K$ is polynomially convex. Since $dp_{1}\wedge\cdots\wedge dp_{n}(z)\neq 0\text{ on }\Gamma,$ $\Gamma$ is totally real, and being graph over a totally real manifold, $K$ is also totally real. Therefore, we get that $\mathscr{P}(K)=\mathcal{C}(K).$ Since $K$ is the graph of $(\overline{p_{1}},\cdots,\overline{p_{n}})$ on $\Gamma,$ hence
$[z_{1},\cdots,z_{n},\overline{p_{1}},\cdots,\overline{p_{n}};\Gamma]=\mathcal{C}(\Gamma).$ (2.14)
Now we will prove the main result. Assume $[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]=\mathcal{C}(\Gamma).$ Define $X={\sf Gr}_{f}(\Gamma).$ Since $[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]=\mathcal{C}(\Gamma),$ then $\mathscr{P}(X)=\mathcal{C}(X).$ Hence $X$ is polynomially convex. Conversely, assume that ${\sf Gr}_{f}(\Gamma)$ is polynomially convex. Let $\Omega$ be a neighborhood of $\Gamma$ such that $p_{1}(z)\not=0~{}~{}\text{ on }\Omega.$ Define $\phi(z,w)=\frac{1}{p_{1}(z)}.$ Then $\phi$ is holomorphic on $\Omega.$ Also, $\phi$ is holomorphic on $\Omega\times\mathbb{C}^{n}.$ Since ${\sf Gr}_{f}(\Gamma)\subset\Omega\times\mathbb{C}^{n},$ by the Oka-Weil approximation theorem, there exists a sequence of polynomials $P_{j}$ in $\mathbb{C}^{n}_{z}\times\mathbb{C}^{n}_{w}$ such that $P_{j}\to\phi$ uniformly on ${\sf Gr}_{f}(\Gamma).$ This implies $P_{j}(z,f(z))\to\phi(z,w)=\frac{1}{p_{1}(z)}=\overline{p_{1}(z)}$ uniformly on $\Gamma.$ Hence $\overline{p_{1}(z)}\in[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma].$ By the similar method we can show that $\overline{{p_{j}}}\in[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]~{}\forall_{j}\in\{1,\cdots,n\}.$ Hence, $[z_{1},\cdots,z_{n},\overline{p_{1}},\cdots,\overline{p_{n}};{\Gamma}]\subset[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma].$ Using (2.14), we get that
$[z_{1},\cdots,z_{n},\overline{p_{1}},\cdots,\overline{p_{n}};\Gamma]=\mathcal{C}(\Gamma)=[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma].$
3. Proof of Theorem 1.2
Let $h_{j}$ denote the pluriharmonic extension of $f_{j}$ to $\mathfrak{D}_{2},$ and write $h=(h_{1},\cdots,h_{N}):\overline{\mathfrak{D}}_{2}\to\mathbb{C}^{N}.$ The following two cases may occur: Case I: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is polynomially convex. In view of Theorem 2.24, we have that
$[z_{1},z_{2},f_{1},\cdots,f_{N};\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$
Therefore, in this case, the proof of the theorem is complete. Case II: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex. By Lemma 2.22, it is enough to find an irreducible analytic variety $V\subset\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ Since ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, there exists $\alpha\in\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Now, we discuss different situations through some lemmas.
Lemma 3.1.
If there exists $\alpha\in\partial{\mathfrak{D}_{2}}-\Gamma_{\mathfrak{D}_{2}}$ such that ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})},$ then there exists an analytic set $\triangle\subset\left(\partial{\mathfrak{D}_{2}}-\Gamma_{\mathfrak{D}_{2}}\right)$ where all $h_{j}$’s are holomorphic.
Proof.
Since $\alpha\in\partial{\mathfrak{D}_{2}}-\Gamma_{\mathfrak{D}_{2}},$ without loss of generality, we assume that $\alpha\in\{z\in\mathbb{C}^{2}:|p_{1}(z)|=1,|p_{2}(z)|<1\}.$ We now consider the set
$\triangle_{\alpha}$ :=$:=\{(z_{1},z_{2})\in\mathfrak{D}_{2}:p_{1}(z_{1},z_{2})=p_{1}(\alpha),|p_{2}(z_{1},z_{2})|<1\}\subset\partial{\mathfrak{D}_{2}},$
$\partial\triangle_{\alpha}$ =$=\{(z_{1},z_{2})\in\mathfrak{D}_{2}:p_{1}(z_{1},z_{2})=p_{1}(\alpha),|p_{2}(z_{1},z_{2})|=1\}.$
We claim that $(\alpha,h(\alpha))\in\widehat{\\ Gr_{h}(\partial\triangle_{\alpha})}:$ if not, then there exists a polynomial $\rho(z,w)$ in $\mathbb{C}^{2+N}$ such that $\rho(\alpha,h(\alpha))=1$ and $\sup_{{\sf Gr}_{h}({\partial\triangle_{\alpha}})}|\rho(z,w)|<1,$ i.e. $\sup_{\partial\triangle_{\alpha}}\left|\rho(z,h(z))\right|<1.$ Next, we consider the function $\psi(z_{1},z_{2})=\frac{1}{2}\left(1+p_{1}(z_{1},z_{2})\overline{p_{1}(\alpha)}\right).$ Note that
$\psi$ $=1\text{ on }{\partial\triangle_{\alpha}}$
$|\psi|$ $<1\text{ on }\Gamma_{\mathfrak{D}_{2}}\setminus{\partial\triangle_{\alpha}}.$
In fact,
$\psi(z_{1},z_{2})$ $=\frac{1}{2}\left(1+p_{1}(z_{1},z_{2})\overline{p_{1}(\alpha)}\right)=\frac{1}{2}\left(p_{1}(\alpha)\overline{p_{1}(\alpha)}+p_{1}(z_{1},z_{2})\overline{p_{1}(\alpha)}\right).$
This implies
$|\psi(z_{1},z_{2})|$ $=\frac{1}{2}\left|p_{1}(\alpha)+p_{1}(z_{1},z_{2})\right|.$
Let $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}).$ We will find $\max|\omega|$ on $\Gamma_{\mathfrak{D}_{2}},$ i.e., we need to find $\max|\omega|$ on $|\omega-p_{1}(\alpha)|=|p_{1}(z)|=1.$ Let $\omega=x+iy~{}~{}\text{ and }p_{1}(\alpha)=\lambda_{1}+i\lambda_{2}.$ The equation of the straight line through the origin and through the centre of the circle $|\omega-p_{1}(\alpha)|=1$ is given by $y=\frac{\lambda_{2}}{\lambda_{1}}x.$ Solving $y=\frac{\lambda_{2}}{\lambda_{1}}x$ and $(x-\lambda_{1})^{2}+(y-\lambda_{2})^{2}=1,$ we get that $(x,y)=(0,0),~{}~{}\text{ or }(x,y)=(2\lambda_{1},2\lambda_{2}).$ Therefore, $\max|\omega|=2$ and maximum attains at $\omega=2\lambda_{1}+i2\lambda_{2}=2p_{1}(\alpha).$ Then $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}),$ implies $p_{1}(z)=p_{1}(\alpha).$ Therefore, $\max|\psi|=1$ on $\Gamma_{\mathfrak{D}_{2}}$ and since $|\omega|<1$ if $\omega\not=2p_{1}(\alpha),$ $|\psi|<1\text{ on }\Gamma_{\mathfrak{D}_{2}}\setminus\partial\triangle_{\alpha}.$ Finally we define the polynomial in $\mathbb{C}^{2+N}$
$\chi(z,w):=\psi^{m}(z)\times\rho(z,w).$
Then
$\chi(\alpha,h(\alpha))$ $=1~{}~{}~{}\text{ and }$
$\sup_{{\sf Gr}_{h}\left(\Gamma_{\mathfrak{D}_{2}}\right)}|\chi(z,w)|$ $=\left|\chi(\eta,h(\eta))\right|\text{ for some }\eta\in\Gamma_{\mathfrak{D}_{2}}.$
If $\eta\in{\partial\triangle_{\alpha}},$ then $\left|\chi(\eta,h(\eta))\right|<1.$ If $\eta\notin{\partial\triangle_{\alpha}},$ since $|\psi|<1$ on $\Gamma_{\mathfrak{D}_{2}}\setminus{\partial\triangle_{\alpha}},$ for sufficiently large $m$ we can make $\left|\chi(\eta,h(\eta))\right|<1.$ Therefore, $(\alpha,h(\alpha))\notin\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ This is a contradiction. Hence, our claim is true. This implies ${\sf Gr}_{h}(\partial\triangle_{\alpha})$ is not polynomially convex. Note that $\{z\in\mathbb{C}^{2}:p_{1}(z)=1\}$ is equivalent to $\mathbb{C}$ (by Result 2.8), and since $\overline{\triangle_{\alpha}}$ is polynomially convex, each component of $\triangle_{\alpha}$ is simply connected. Also, by using Lemma 2.15, we have that $h_{j}$’s are harmonic on $\triangle_{\alpha}.$ Then by applying the Wermer maximality theorem, we obtain that all $h_{j}$’s are holomorphic on some component of $\triangle_{a}$ and ${\sf Gr}_{h}(\triangle_{\alpha})\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$
From next, we assume that
$\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap{\sf Gr}_{h}\left(\partial{\mathfrak{D}_{2}}-\Gamma_{\mathfrak{D}_{2}}\right)=\emptyset.$ (3.1)
We set
$\widetilde{V}:=\{z\in\mathfrak{D}_{2}:\bar{\partial}{{h_{j_{1}}}}\wedge\bar{\partial}{{h_{j_{2}}}}(z)=0,\forall~{}{j_{1},j_{2}},1\leq j_{1},j_{2}\leq N\}.$ (3.2)
Because of Result 2.10 and using the fact that the finite intersection of analytic varieties is again an analytic variety, we get that $\widetilde{V}$ is an analytic variety in $\mathfrak{D}_{2}$. By Result 2.1, we get that ${\sf Gr}_{h}(\mathfrak{D}_{2}\setminus\widetilde{V})$ is totally real. Let $E={\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ and $\Omega=(\mathfrak{D}_{2}\setminus\widetilde{V})\times\mathbb{C}^{N}.$ Then
$\widehat{E}\cap\Omega\subset\left({\sf Gr}_{h}(\overline{\mathfrak{D}}_{2})\cap(\mathfrak{D}_{2}\setminus\widetilde{V})\times\mathbb{C}^{N}\right)\subset{\sf Gr}_{h}(\mathfrak{D}_{2}\setminus\widetilde{V}).$
Since ${\sf Gr}_{h}(\mathfrak{D}_{2}\setminus\widetilde{V})$ is totally real, by Lemma 2.23, we get that $\widehat{E}\cap\Omega=\emptyset.$ Hence
$\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap{\sf Gr}_{h}(\mathfrak{D}_{2}\setminus\widetilde{V})\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap(\mathfrak{D}_{2}\setminus\widetilde{V})\times\mathbb{C}^{N}=\emptyset.$ (3.3)
In view of (3.1) and (3.3), we get that
$\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}}\cup\widetilde{V}),$
i.e.
$\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\setminus{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})\subset{\sf Gr}_{h}(\widetilde{V}).$ (3.4)
Note that $\widetilde{V}$ can not be discrete. Therefore, two cases can occur: either $\widetilde{V}\neq\mathfrak{D}_{2}$ or $\widetilde{V}=\mathfrak{D}_{2}.$ First, we consider the case $\widetilde{V}\neq\mathfrak{D}_{2}.$
Lemma 3.2.
[…]
Proof.
We choose a neighborhood $U$ of $\alpha$ in $V$ such that $U\cap V=:D_{\alpha}$ is a smoothly bounded disc containing $\alpha.$ Let us define $X:=\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial U\times\mathbb{C}^{N}).$ Note that $X$ is a closed subset of a compact set $\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})},$ and hence $X$ is compact. Since $\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}}\cup\widetilde{V}),$ we have
$X\subset{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}}\cup\widetilde{V})\cap(\partial U\times\mathbb{C}^{N})\subset{\sf Gr}_{h}(\partial U\cap V).$
We claim that
${\sf Gr}_{h}(\alpha)\in\widehat{X}\subset\widehat{{\sf Gr}_{h}(\partial U\cap V)}=\widehat{{\sf Gr}_{h}(\partial D_{\alpha})}.$
To prove this claim, let $E={\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ and $N=U\times\mathbb{B}(r)$ with $r>\sup_{\overline{U}}|h|.$ Note that $E\cap N=\emptyset$ and $\partial N\cap\widehat{E}=[(\partial U\times\mathbb{B}(r))\cap\widehat{E}]\cup[(U\times\partial\mathbb{B}(r))\cap\widehat{E}].$ We claim that $(U\times\partial\mathbb{B}(r))\cap\widehat{E}=\emptyset:$ if possible, let $(\eta,\xi)\in(U\times\partial\mathbb{B}(r))\cap\widehat{E}.$ Then $\xi=h(\eta)$ and $\eta\in U,|\xi|=r.$ But on $U,$ we have $|\xi|=|h(\eta)|<r.$ This is a contradiction to the assumption that $(U\times\partial\mathbb{B}(r))\cap\widehat{E}\not=\emptyset.$ Then by Result 2.4, we have
${\sf Gr}_{h}(\alpha)\in\widehat{\widehat{E}\cap\partial N}$ $=\widehat{[\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial U\times\mathbb{B}(r))]}$
$\subset\widehat{[\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial U\times\mathbb{C}^{N})]}=\widehat{X}\subset\widehat{{\sf Gr}_{h}(\partial U\cap V)}=\widehat{{\sf Gr}_{h}(\partial D_{\alpha})}.$
Therefore, $\widehat{{\sf Gr}_{h}(\partial D_{\alpha})}$ is not polynomially convex, and hence by the Wermer maximality theorem we can say that all $h_{j}$’s are holomorphic in $D_{\alpha}.$ Note that $X={\sf Gr}_{h}(\partial D_{\alpha}).$ In fact, if $X$ is proper subset of ${\sf Gr}_{h}(\partial D_{\alpha}),$ then $X={\sf Gr}_{h}(K)$ for some proper compact subset of $\partial D_{\alpha}$ and in this case $X$ will be polynomially convex. Hence ${\sf Gr}_{h}(\alpha)\in\widehat{X}.$ This is a contradiction and hence $X={\sf Gr}_{h}(\partial D_{\alpha}).$ It is easy to see that ${\sf Gr}_{h}(\overline{D_{\alpha}})\subset\widehat{X}.$ To see this, let $(a,h(a))\in{\sf Gr}_{h}(\overline{D_{\alpha}}),$ and $P$ be any holomorphic polynomial in $\mathbb{C}^{2}\times\mathbb{C}^{N}.$ Then
$|P((a,h(a)))|\leq$ $\sup_{{\sf Gr}_{h}(\overline{D_{\alpha}})}|P(z,w)|$
$=$ $\sup_{(\partial D_{\alpha})}|P(z,h(z))|~{}(\text{since }h\text{ is holomorphic on}~{}D_{\alpha}).$
Hence $(a,h(a))\in\widehat{X}.$ Therefore, ${\sf Gr}_{h}(D_{\alpha})\subset\widehat{X}\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}$ and this holds near every regular point. Since regular points are dense, it follows that if $\beta$ is a singular point of $V,$ we have ${\sf Gr}_{h}(\beta)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}$ also true. Hence ${\sf Gr}_{h}(V)\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})},$ and in view of (3.1), we have that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$
From now, we assume that
$\widetilde{V}=\mathfrak{D}_{2}.$
In view of Lemma 2.14, we can say that there exist holomorphic functions $\Phi_{j}$ and $\Psi_{j}$ on $\mathfrak{D}_{2}$ such that
${\sf Im}h_{j}={\sf Re}\Psi_{j}\text{ and }u_{j}:={\sf Re}\Phi_{j}=h_{j}-\Psi_{j}.$ (3.5)
We use the following change of coordinates $\widetilde{\Psi}:\mathfrak{D}_{2}\times\mathbb{C}^{N}\mapsto\mathfrak{D}_{2}\times\mathbb{C}^{N}$ on $\mathfrak{D}_{2}\times\mathbb{C}^{N}$ defined by
$\widetilde{\Psi}(z_{1},z_{2},w_{1},\cdots,w_{N})=(z_{1},z_{2},w_{1}-\Psi_{1}(z),\cdots,w_{N}-\Psi_{N}(z)).$
We set $u=(u_{1},u_{2},\cdots,u_{N})$ and $\Psi(z)=\left(\Psi_{1}(z),\cdots,\Psi_{N}(z)\right).$ Let $K$ be any compact subset of $\mathfrak{D}_{2}.$ Then $\widetilde{\Psi}\left(\widehat{{\sf Gr}_{h}}(K)\right)=\widehat{{\sf Gr}_{u}(K)}.$ Therefore, we have the following:
Lemma 3.3.
Let $K$ be any compact subset of $\mathfrak{D}_{2}.$ Then $\widetilde{\Psi}\left(\widehat{{\sf Gr}_{h}}(K)\right)=\widehat{{\sf Gr}_{u}(K)}.$ Therefore, a point ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(K)}$ is equivalent to ${\sf Gr}_{u}(\alpha)\in\widehat{{\sf Gr}_{u}(K)}.$
By (3.1) and the fact that $\widetilde{V}=\mathfrak{D}_{2},$ we get that
$\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{h}\left(\Gamma_{\mathfrak{D}_{2}}\right)\cup{\sf Gr}_{h}\left({\mathfrak{D}_{2}}\right).$
Therefore $\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{h}(D),$ for some subset $D$ of $\overline{\mathfrak{D}_{2}}.$ Since polynomial convex hull of a compact set is compact, $D$ is also compact. Let $r_{j}$ be an increasing sequence of real numbers such that $r_{j}\to 1$ as $j\to\infty.$ We define
$R^{1}_{j}$ $:=\min\left\{|p_{1}(z)|:\{|p_{1}(z)|\leq r_{j},|p_{2}(z)|=r_{j}\}\cap D\neq\emptyset\right\}$
$R^{2}_{j}$ $:=\min\left\{|p_{2}(z)|:\{|p_{2}(z)|\leq r_{j},|p_{1}(z)|=r_{j}\}\cap D\neq\emptyset\right\}.$
Observe that $R^{1}_{j}=p_{1}(\xi_{j})$ for some $\xi_{j}\in D,$ and $\xi_{j}\to\Gamma_{\mathfrak{D}_{2}}$ as $j\to\infty.$ Therefore, $j\to\infty$ implies $R^{1}_{j}\to 1.$ Similarly, $R^{2}_{j}\to 1$ as $j\to\infty.$ Define a sequence of real numbers
$\varepsilon_{j}:=\max\left\{r_{j}-R^{1}_{j},r_{j}-R^{2}_{j}\right\}.$
Since $r_{j}-R^{1}_{j},r_{j}-R^{2}_{j}\to 0$ as $j\to\infty,$ therefore, $\varepsilon_{j}\to 0$ as $j\to\infty.$ Define
$\mathfrak{Q}_{j}:=\{z\in\mathfrak{D}_{2}:|p_{1}|=r_{j},r_{j}-\varepsilon_{j}\leq|p_{2}|\leq r_{j}\}\cup\{z\in\mathfrak{D}_{2}:|p_{2}|=r_{j},r_{j}-\varepsilon_{j}\leq|p_{1}|\leq r_{j}\}.$
and
$\overline{\mathfrak{D}^{r_{j}}_{2}}=\{z\in\mathfrak{D}_{2}:|p_{2}|\leq r_{j},|p_{1}|\leq r_{j}\}.$
Then from the above construction, we obtained that
$\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap{\sf Gr}_{h}(\partial\mathfrak{D}^{r_{j}}_{2})\subset{\sf Gr}_{h}(\mathfrak{Q}_{j}).$ (3.6)
Lemma 3.4.
Assume that there is a subsequence $r_{j_{k}}$ of $r_{j}$ such that for each $k$ there exists a point ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j_{k}})}$ with $|p_{1}(a_{k},b_{k})|=r_{j_{k}}$ and $|p_{2}(a_{k},b_{k})|<r_{j_{k}}-\varepsilon_{j_{k}}.$ Then there exists a analytic set normal-$\triangle$ and all $h_{j}$’s are holomorphic on .normal-$\triangle$.△ .
Proof.
Let $\xi\in\{|p_{1}|(z)=r_{j_{k}},r_{j_{k}}-\varepsilon_{j_{k}}\leq|p_{2}(z)|\leq r_{j_{k}}\}\subset\mathfrak{Q}_{j_{k}}.$ Then $|p_{1}(\xi)|=r_{j_{k}}.$ Also $|p_{1}(a_{k},b_{k})|=r_{j_{k}}.$ Therefore, $p_{1}(\xi)=e^{i\theta_{0}}p_{1}(a_{k},b_{k})$ for some $\theta_{0}.$ For each $k,$ we consider the set
$\triangle_{k}:=\{z:p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),r_{j_{k}}-\varepsilon_{j_{k}}\leq|p_{2}(z)|\leq r_{j_{k}}\}.$
We claim that ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\triangle_{k})}:$ if not, then there exists a polynomial ϝ$\digamma_{k}$ such that
ϝ$\digamma_{k}((a_{k},b_{k}),h(a_{k},b_{k}))$ $=1\text{ and }\sup_{{\sf Gr}_{h}(\triangle_{k})}|\digamma_{k}(z,w)|<1.$
This implies
$\sup_{\triangle_{k}}|\digamma_{k}(z,h(z))|$ $<1.$ (3.7)
Again, consider the polynomial
$\Upsilon_{k}(z_{1},z_{2}):=\frac{1}{2}\left[1+\frac{p_{1}(z_{1},z_{2})e^{-i\theta_{0}}\overline{p_{1}(a_{k},b_{k})}}{r^{2}_{j_{k}}}\right].$ (3.8)
We observe that $\Upsilon_{k}=1\text{ if }p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k})$ and
[…]
Finally, consider the polynomial
$\kappa_{k}(z,w)=\digamma_{k}(z,w)\Upsilon^{m}_{k}(z)$
Hence $\kappa_{k}\left((a_{k},b_{k}),h(a_{k},b_{k})\right)=1.$ We claim that
$\sup_{{\sf Gr}_{h}(\mathfrak{Q}_{j})}|\kappa_{k}(z,w)|.$
To see a proof of the claim, let $\sup_{{\sf Gr}_{h}(\mathfrak{Q}_{j})}|\kappa_{k}(z,w)|=|\kappa_{k}(s,h(s))|,$ for some $s=(s_{1},s_{2})\in\mathfrak{Q}_{j}.$ If $s\in\triangle_{k},$ then from (3.7) and (3.8), we get that $|\kappa_{k}(s,h(s))|<1.$ If $s\in\mathfrak{Q}_{j}\setminus\triangle_{k},$ then for large enough $m,$ we have that $|\kappa_{k}(s,h(s))|<1.$ This is a contradiction. Hence ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\triangle_{k})}.$ This implies ${\sf Gr}_{u}(a_{k},b_{k})\in\widehat{{\sf Gr}_{u}(\triangle_{k})},$ where $u$ is given by 3.5. Note that $\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k})\}$ is equivalent to $\mathbb{C}$ (by Result 2.8), and $u_{j}$’s are harmonic on $\triangle_{k}.$ Hence $u$ is constant on some component of $\triangle_{k},$ i.e. all $h_{j}$’s are holomorphic on that component of $\triangle_{k}.$ We will denote this component again by $\triangle_{k}.$ We now claim that ${\sf Gr}_{h}(\triangle_{k})\subset\widehat{{\sf Gr}_{h}(\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\})}.$ To prove this claim, let $(\xi,h(\xi))\in{\sf Gr}_{h}(\triangle_{k})$ for some $\xi\in\triangle_{k}.$ Then for any polynomial $P(z,w)$ in $\mathbb{C}^{2},$
$|P(\xi,h(\xi))|$ $\leq\sup_{{\sf Gr}_{h}(\triangle_{k})}|P(z,w)|$
$=\sup_{\triangle_{k}}|P(z,h(z))|$
$=\sup\{|P(z,h(z))|:\;p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\}.$
The last equality holds because $h$ is holomorphic on $\triangle_{k}$. Therefore,
$(\xi,h(\xi))\in\widehat{{\sf Gr}_{h}(\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\})}.$
This implies that ${\sf Gr}_{h}(\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\})$ is not polynomially convex. Hence, $h$ is holomorphic on some simply connected component $\widetilde{\triangle_{k}}$ of $\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|\leq r_{j_{k}}\}.$ That is, there exists biholomorphism $\phi_{k}:\mathbb{D}\to\widetilde{\triangle_{k}}$ such that $h_{j}\circ\phi_{k}:\mathbb{D}\to\mathbb{C}$ is holomorphic $\forall j=1,\cdots,N;~{}\forall k\geq 1.$ Again there exists a subsequence of $(a_{k},b_{k})$ which is convergent to some point $(a,b).$ Since $|p_{1}(a_{k},b_{k})|=r_{j_{k}}$ convergent to $1$ (for the above subsequence), then $p_{1}{(a_{k},b_{k})}$ converges to $p_{1}(a,b)$ and $|p_{1}(a,b)|=1.$ Hence $\phi_{k}(\mathbb{D})=\widetilde{\triangle_{k}}$ converges to some simply connected component $\widetilde{\triangle}$ of $\{p_{1}(z)=p_{1}(a,b),|p_{2}(z)|\leq 1\}.$ Therefore, there exists a biholomorphism $\phi:\mathbb{D}\to\widetilde{\triangle}$ such that $h_{j}\circ\phi:\mathbb{D}\to\mathbb{C}$ is holomorphic $\forall j=1,\cdots,N.$
Without loss of generality, we assume that $u_{1}$ is non-constant. For $c\in\mathbb{R},$ define $L_{c}=L_{c}(u_{1}):=\{z\in\mathfrak{D}_{2}:u_{1}(z)=c\}.$ Since ${\sf Re}\Phi_{1}=u_{1},$ we have
$L_{c}(u_{1})=\cup_{r\in Img(\Phi_{1})}L_{c+i.r}(\Phi_{1})$
is the disjoint union of analytic (since $\Phi_{1}$ is holomorphic) sets, each of them is leaf of $L_{c}.$ If $\Phi_{1}(z_{0})=c+ir,$ for some $c+ir\in Img(\Phi_{1}),$ then we denote $\mathfrak{L}_{z_{0}}:=\{z\in\mathfrak{D}_{2}:\Phi_{1}(z)=\Phi_{1}(z_{0})\}$ is the leaf through $z_{0}.$
Lemma 3.5.
There exists a discrete set $W\subset\mathfrak{D}_{2}$ such that near any point $z_{0}\in\mathfrak{D}_{2}\setminus W,$ the set $\{\Phi_{1}(z)=\Phi_{1}(z_{0})\}$ is a smooth surface.
Proof.
Let $Y:=\{\Phi_{1}(z)=\Phi_{1}(z_{0})\}$ be a complex analytic variety of $\mathfrak{D}_{2}.$ Let $W$ be the collection of singular points of $Y.$ Being a singular set of a complex analytic variety $Y,$ $W$ is also a complex analytic variety with $\dim_{\mathbb{C}}W<\dim_{\mathbb{C}}Y=1.$ Hence $W$ is discrete. Therefore $Y\setminus W$ form a complex manifold of dimension one (by Implicit function theorem). ∎
Lemma 3.6.
Assume that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}\cap{\sf Gr}_{h}(\mathfrak{D}^{r_{j}}_{2})$ and set $c_{1}=u_{1}(\alpha_{0}).$ Then there exists a point $z_{0}\in(L_{c_{1}}\setminus W)\cap\mathfrak{D}^{r_{j}}_{2}$ with ${\sf Gr}_{h}(z_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}\cap{\sf Gr}_{h}(\mathfrak{D}^{r_{j}}_{2}).$
Proof.
Since $\mathfrak{Q}_{j}$ is a compact subset of $\mathfrak{D}_{2}$ and ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}$, by Lemma 3.3, ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}.$ Let us define
${\bf c}:=(c_{1},\cdots,c_{N})=u(\alpha_{0}),\text{ and }\quad\mathfrak{Q}^{{\bf c}}_{j}:=\{z\in\mathfrak{Q}_{j}:u(z)={\bf c}\}.$
We consider the map $F:\mathbb{C}_{z}^{2}\times\mathbb{C}_{w}^{N}\to\mathbb{C}^{N}$ defined by
$F(z,w)=w-{\bf c}.$
Let $K:={\sf Gr}_{u}(\mathfrak{Q}_{j}).$ Then,
  • $Y:=F(K)=\{u(z)-{\bf c}:z\in\mathfrak{Q}_{j}\}\subset\mathbb{R}^{N}$;
  • $0\in Y,$ because $\alpha_{0}\in\mathfrak{Q}^{{\bf c}}_{j}\subset\mathfrak{Q}_{j}$;
  • $F^{-1}\{0\}\cap K={\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})$: let
    $(z,w)\in F^{-1}\{0\}\cap K$ $\Longleftrightarrow F(z,w)=0\text{ and }(z,w)\in{\sf Gr}_{u}(\mathfrak{Q}_{j})$
    $\Longleftrightarrow w={\bf c}\text{ and }(w=u(z),z\in\mathfrak{Q}_{j})$
    $\Longleftrightarrow(z,w)\in{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j}).$
Therefore, by Result 2.5, we obtain that
$\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}\cap F^{-1}(0)$ $=\widehat{({\sf Gr}_{u}(\mathfrak{Q}_{j})\cap F^{-1}\{0\}})=\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})}.$
Since ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}$ and ${\sf Gr}_{u}(\alpha_{0})\in F^{-1}\{0\},$ hence ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})}.$ Now,
$\mathfrak{Q}^{{\bf c}}_{j}\subset\widetilde{L_{{\bf c}}}:=\{z\in\overline{\mathfrak{D}^{r_{j}}_{2}}:u(z)={\bf c}\}$
and $\widetilde{L_{{\bf c}}}$ is polynomially convex. Therefore,
$\widehat{\mathfrak{Q}^{\bf c}_{j}}\subset\widehat{\widetilde{L_{\bf c}}}=\widetilde{L_{\bf c}}\subset L_{c_{1}}.$
This implies
$\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{\bf c}_{j})}\subset\widehat{{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})}\subset\widehat{\mathfrak{Q}^{\bf c}_{j}}\times\mathbb{C}^{N}\subset L_{c_{1}}\times\mathbb{C}^{N}.$
Hence, ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{\bf c}_{j})}$ implies
$\alpha_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}\subset\widetilde{L_{\bf c}}\subset L_{c_{1}}.$ (3.9)
If $\alpha_{0}\in L_{c_{1}}\setminus W,$ we are done. If $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains only singular points and since singular points are isolated, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ is finite. Therefore, $\alpha_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}=\mathfrak{Q}^{\bf c}_{j},$ but we have taken $\alpha_{0}$ from $\mathfrak{D}^{r_{j}}_{2}.$ This is a contradiction. Therefore, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains at least one point $z_{0}\in L_{c_{1}}\setminus W.$ We claim that ${\sf Gr}_{u}(z_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}.$ To prove this claim, let $H(z,w)$ be a polynomial in $\mathbb{C}^{2+N}.$ Since $(z_{0},u(z_{0}))\in{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})$, then
$|H(z_{0},u(z_{0}))|$ $\leq\sup_{{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})}|H(z,w)|$
$=\sup_{\widehat{\mathfrak{Q}^{\bf c}_{j}}}|H(z,u(z))|$
$=\sup_{\widehat{\mathfrak{Q}^{\bf c}_{j}}}|H(z,\bf c)|,\text{ since }\widehat{\mathfrak{Q}^{\bf c}_{j}}\subset\widetilde{L_{\bf c}}$
$\leq\sup_{\mathfrak{Q}^{\bf c}_{j}}|H(z,\bf c)|$
$=\sup_{\mathfrak{Q}^{\bf c}_{j}}|H(z,u(z))|,\text{ since }\mathfrak{Q}^{\bf c}_{j}\subset\widetilde{L_{\bf c}}$
$\leq\sup_{\mathfrak{Q}_{j}}|H(z,u(z))|,\text{ since }\mathfrak{Q}^{\bf c}_{j}\subset\mathfrak{Q}_{j}.$
This implies $(z_{0},u(z_{0}))\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}.$ In view Lemma 3.3, we have that $(z,h(z))\in{\sf Gr}_{h}(\widehat{\mathfrak{Q}_{j}})$ which completes the proof. ∎
Lemma 3.7.
For any point $z_{0}\in L_{c_{1}}\setminus W,$ we have that ${\sf Gr}_{h}(z_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}$ implies that ${\sf Gr}_{h}(\mathfrak{L}_{z_{0}}\cap\mathfrak{D}^{r_{j}}_{2})\subset\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}$ with $h$ holomorphic along $\mathfrak{L}_{z_{0}}$ where $\mathfrak{L}_{z_{0}}$ denotes the leaf through $z_{0}.$
First, we prove the following result.
Lemma 3.8.
Let $z_{0}\in(L_{c_{1}}\setminus W)\cap\mathfrak{D}^{r_{j}}_{2}$ and $\mathfrak{L}_{z_{0}}$ be the leaf through $z_{0}.$ Then $z_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}$ if and only if $\mathfrak{L}_{z_{0}}\subset\widehat{\mathfrak{Q}^{\bf c}_{j}}.$
Proof.
If $\mathfrak{L}_{z_{0}}\subset\widehat{\mathfrak{Q}^{\bf c}_{j}},$ then $z_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}.$ We need to show only the other side. Let $U$ be a neighborhood of $z_{0}$ and $D(z_{0})=\mathfrak{L}_{z_{0}}\cap U$ is a smoothly bounded disc containing $z_{0}.$ Using a translation by $-\Phi_{1}(z_{0}),$ without loss of generality, we may assume that $\Phi_{1}(z_{0})=0\implies c_{1}=0.$ For $z\in L_{c_{1}},u_{1}(z)+iv_{1}(z)=iv_{1}(z)\in i\mathbb{R}.$ Hence
$\Phi_{1}(L_{c_{1}})\subset i\mathbb{R}\text{ and }\Phi_{1}(L_{c_{1}}\setminus\mathfrak{L}_{z_{0}})\subset i\mathbb{R}\setminus\{0\}.$ (3.10)
Let $X:=\widehat{\mathfrak{Q}^{\bf c}_{j}}\cap\partial U.$ By Result 2.4, we get that $z_{0}\in\widehat{X}.$ We claim that $\partial D(z_{0})=(\partial U\cap\mathfrak{L}_{z_{0}})\subset X.$ If not, consider the set $X_{z_{0}}:=\partial D(z_{0})\cap X.$ Then $X_{z_{0}}$ is a proper subset of $\partial D(z_{0}).$ This implies $z_{0}\notin\widehat{X_{z_{0}}}.$ Then there exists a polynomial $\gamma$ such that
$\gamma(z_{0})=1~{}~{}~{}\text{ and }\sup_{X_{z_{0}}}|\gamma|<1.$
Again, consider the function $\xi(t)=e^{t^{2}}$ from $\mathbb{C}$ to $\mathbb{C}.$ We consider the function
$\phi(z_{1},z_{2})=\gamma(z_{1},z_{2})\times(\xi\circ\Phi_{1})^{d}(z_{1},z_{2}),~{}d\in\mathbb{N}.$
Then $\phi(z_{0})=\gamma(z_{0})\times(\xi\circ\Phi_{1}(z_{0}))^{d}=1.$ Now we show that $\sup_{X}|\phi|<1.$ Since $X$ is compact, $\sup_{X}|\phi|=|\phi(s)|$ for some $s\in X.$ Note that
[…]
$\subset(L_{\bf c}\cap\bar{U}\setminus D(z_{0}))\cup(\mathfrak{L}_{z_{0}}\cap\partial U).$
Case I If $s\in\mathfrak{L}_{z_{0}}\cap\partial U$ and since $s\in X,$ therefore $s\in X_{z_{0}}=X\cap\partial D(z_{0}),$ then we have $|\gamma(s)|<1.$ Also, since $\Phi_{1}(s)\in i\mathbb{R},$ there exists $y\in\mathbb{R}$ such that $\Phi_{1}(s)=iy.$ Therefore, $(\xi\circ\Phi_{1}(s))=e^{-y^{2}}<1.$ This implies $|(\xi\circ\Phi_{1}(s))^{d}|\leq 1.$ This contradicts that $z_{0}\in\widehat{X}.$ Case II If $s\in(L_{\bf c}\cap\bar{U}\setminus D(z_{0}))$ and $s\notin\mathfrak{L}_{z_{0}}\cap\partial U$ implies $s\notin\mathfrak{L}_{z_{0}},$ by (3.10), $|\xi\circ\Phi_{1}(s))|<1.$ Then for sufficiently large $d$, we have $|\gamma(s)|\times|\xi\circ\Phi_{1}(s))|^{d}<1.$ Again this is a contradiction to the fact that $z_{0}\in\widehat{X}.$ Therefore, $X_{z_{0}}=\partial D(z_{0})\subset X.$ Hence,
$D(z_{0})\subset\widehat{\partial D(z_{0})}\subset\widehat{X}\subset\widehat{\mathfrak{Q}^{\bf c}_{j}}\subset\widetilde{L_{\bf c}}.$ (3.11)
Proof of Lemma 3.8 follows as $\mathfrak{L}_{z_{0}}$ can be covered by such $D(z_{0}).$
We now complete the proof of Lemma 3.7.
Proof of Lemma 3.6.
let $z_{0}$ be as in the hypothesis. Then by Lemma 3.6, $z_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}.$ By Lemma 3.8, $\mathfrak{L}_{z_{0}}\subset\widehat{\mathfrak{Q}^{\bf c}_{j}}.$ For $z\in\mathfrak{L}_{z_{0}}\cap\mathfrak{D}^{r_{j}}_{2},$ and for any polynomial $Q(z,w)$ we have:
$|Q(z,u(z))|$ $\leq\sup_{{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})}|Q(z,w)|$
$=\sup_{\widehat{\mathfrak{Q}^{\bf c}_{j}}}|Q(z,{\bf c})|,\text{ since }\widehat{\mathfrak{Q}^{\bf c}_{j}}\subset\widetilde{L_{\bf c}}$
$=\sup_{\mathfrak{Q}^{\bf c}_{j}}|Q(z,{\bf c})|$
$=\sup_{\mathfrak{Q}^{\bf c}_{j}}|Q(z,u(z))|,\text{ since }\mathfrak{Q}^{\bf c}_{j}\subset\widetilde{L_{\bf c}}$
$\leq\sup_{\mathfrak{Q}_{j}}|Q(z,u(z))|.$
Therefore, $(z,u(z))\in{\sf Gr}_{u}(\widehat{\mathfrak{Q}_{j}}).$ Hence $(z,h(z))\in{\sf Gr}_{h}(\widehat{\mathfrak{Q}_{j}}).$ From (3.11), we get that $u_{2},\cdots,u_{N}$ are also constant on $D(z_{0}).$ Hence $h_{j}$’s are holomorphic on $D(z_{0})=U\cap\mathfrak{L}_{z_{0}}$ which completes the proof of Lemma 3.7. ∎
Proof of Theorem 1.2.
If there exists an analytic set $\triangle\subset\left(\partial{\mathfrak{D}_{2}}-\Gamma_{\mathfrak{D}_{2}}\right),$ where all $h_{j}$’s are holomorphic, then there is nothing to prove. We assume that no such analytic set exists where all $h_{j}$’s are holomorphic. Therefore, by Lemma 3.1, we have $\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap{\sf Gr}_{h}\left(\partial{\mathfrak{D}_{2}}-\Gamma_{\mathfrak{D}_{2}}\right)=\emptyset.$ Recall the definition of $\widetilde{V}$ (see 3.2):
$\widetilde{V}:=\{z\in\mathfrak{D}_{2}:\bar{\partial}{{h_{j_{1}}}}\wedge\bar{\partial}{{h_{j_{2}}}}(z)=0,\forall~{}{j_{1},j_{2}},1\leq j_{1},j_{2}\leq N\}.$
Since, by assumption in the beginning of the proof, ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, by (3.4), there exist $\alpha_{0}\in\widetilde{V}$ such that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Case I: Assume that $\widetilde{V}\not=\mathfrak{D}_{2}.$ Then Lemma 3.2 gives an irreducible component $V$ of $\widetilde{V}$ such that all $h_{j}$’s are holomorphic on $V.$ Observe that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ The proof of the theorem is complete in this case in view of Lemma 2.22. Case II: Now we consider the case that $\widetilde{V}=\mathfrak{D}_{2}.$ Without loss of generality, we assume that $h_{1}$ is non-holomorphic. Since ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})},$ there exists $j_{0}\in\mathbb{N}$ such that $\alpha_{0}\in\mathfrak{D}^{r_{j}}_{2}$ for $j>j_{0}.$ Therefore, by (3.6) and Result 2.4, we obtain that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}:$ in fact, assume that
$K:=(\partial\mathfrak{D}^{r_{j}}_{2}\times\mathbb{C}^{N})\cap\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ (3.12)
Since $\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}}\cup\mathfrak{D}_{2}),$ therefore,
$K\subset{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}}\cup\mathfrak{D}_{2})\cap(\partial\mathfrak{D}^{r_{j}}_{2}\times\mathbb{C}^{N})={\sf Gr}_{h}(\partial\mathfrak{D}^{r_{j}}_{2}).$ (3.13)
Next we show that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{K}.$ To show this we assume that $E={\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ and $N=\mathfrak{D}^{r_{j}}_{2}\times\mathbb{B}(r)$ with $r>\sup_{\mathfrak{D}^{r_{j}}_{2}}|h|.$ Note that $E\cap N=\emptyset$ and $\partial N\cap\widehat{E}=(\partial\mathfrak{D}^{r_{j}}_{2}\times\mathbb{B}(r))\cap\widehat{E}\cup(\mathfrak{D}^{r_{j}}_{2}\times\partial\mathbb{B}(r))\cap\widehat{E}.$ Note that $(\mathfrak{D}^{r_{j}}_{2}\times\partial\mathbb{B}(r))\cap\widehat{E}=\emptyset.$ Then, by Result 2.4, we have
${\sf Gr}_{h}(\alpha_{0})\in\widehat{\widehat{E}\cap\partial N}$ $=\widehat{[\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial\mathfrak{D}^{r_{j}}_{2}\times\mathbb{B}(r))]}$
$\subset\widehat{[\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial\mathfrak{D}^{r_{j}}_{2}\times\mathbb{C}^{N})]}=\widehat{K}.$
Using (3.12), (3.13) and(3.6), we obtain that
$K\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}\cap{\sf Gr}_{h}(\partial\mathfrak{D}^{r_{j}}_{2})\subset{\sf Gr}_{h}(\mathfrak{Q}_{j}).$
Therefore, ${\sf Gr}_{h}(\alpha_{0})\in\widehat{K}$ implies
${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}.$
Therefore, we have ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}\cap{\sf Gr}_{h}(\mathfrak{D}^{r_{j}}_{2}).$ By Lemma 3.6, we may assume that $\alpha_{0}\notin W.$ Let $V^{{}^{\prime}}$ denotes the irreducible component of $\{z\in\mathfrak{D}_{2}:\Phi_{1}(z)=\Phi_{1}(\alpha_{0})\}$ containing $\alpha_{0}.$ Then from Lemma 3.7, we get that
${\sf Gr}_{h}(V^{{}^{\prime}}\cap\mathfrak{D}^{r_{j}}_{2})\subset\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j})}\text{ for all }j>j_{0}.$ (3.14)
We now claim that ${\sf Gr}_{h}(V^{{}^{\prime}})\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ To prove this claim, assume that $(\alpha,h(\alpha))\in{\sf Gr}_{h}(V^{{}^{\prime}}).$ Since $\alpha\in V^{{}^{\prime}},$ $\alpha\in V^{{}^{\prime}}\cap\mathfrak{D}^{r_{j_{1}}}_{2}$ for some $j_{1}.$ Therefore, $\alpha\in V^{{}^{\prime}}\cap\mathfrak{D}^{r_{j}}_{2}$ for all $j>j_{1}.$ Using (3.14), for any polynomial $H$ in $\mathbb{C}^{N+2}$ and for $j>\max\{j_{0},j_{1}\},$ we get that
$|H(\alpha,h(\alpha))|\leq\sup_{\mathfrak{Q}_{j}}|H(z,h(z))|=|H(a_{j},h(a_{j}))|$ (3.15)
for some $a_{j}\in\mathfrak{Q}_{j}.$ Taking limit $j\to\infty,$ we obtain that $|H(a_{j},h(a_{j}))|\to H(a,h(a))$ for some $a\in\Gamma_{\mathfrak{D}_{2}}.$ Hence from (3.15), we get that
$|H(\alpha,h(\alpha))|\leq\lim_{j\to\infty}|H(a_{j},h(a_{j}))|=|H(a,h(a))|\leq\sup_{\Gamma_{\mathfrak{D}_{2}}}|H(z,h(z))|.$
Since $H$ is arbitrary, then we can say that $(\alpha,h(\alpha))\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ This proves our claim. Note that $\overline{V^{{}^{\prime}}}\setminus V^{{}^{\prime}}\subset\Gamma_{\mathfrak{D}_{2}}.$ The proof of the theorem is complete in this case in view of Lemma 2.22. ∎
4. Description of hull
We consider the proper polynomial map $\Psi:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\Psi(z)=$ $(p_{1}(z),p_{2}(z)).$ Then the distinguished boundary $\Gamma_{\mathfrak{D}_{2}}$ of the polynomial polyhedron $\mathfrak{D}_{2}=\Psi^{-1}(\mathbb{D}^{2})=\{z\in\mathbb{C}^{2}:|p_{1}(z)|\leq 1,|p_{2}(z)|\leq 1\}$ is $\Gamma_{{\mathfrak{D}_{2}}}=\{z\in\mathbb{C}^{2}:|p_{1}(z)|=1,|p_{2}(z)|=1\}.$ Let $P(z_{1},z_{2})$ be any holomorphic polynomial of degree $m$ in $z_{1}$ and of degree $n$ in $z_{2}.$ We wish to compute $\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}$ under the assumption that $\mathfrak{D}_{2}$ is complex non-degenerate.
First, we state and prove a lemma that will be useful in this work. For a subset $X$ of $\mathbb{C}^{n},$ we denote
#$\#CC(X)=\text{the number of connected components of }X.$
Lemma 4.1.
Let $X$ be a compact subset of $\mathbb{C}^{n}.$ Then #$\#CC(\widehat{X})\leq\#CC(X).$
Proof.
If #$\#CC(X)=\infty,$ then there is nothing to prove. In view Result 2.11, it is not possible that #$\#CC(X)<\infty,$ but #$\#CC(\widehat{X})=\infty.$ In fact, assume that $\widehat{X}=\cup^{\infty}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ This implies $X=X\cap\widehat{X}=\cup^{\infty}_{j=1}X_{j}\cap X.$ Result 2.11 says that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ Since each $X_{j}\cap X$ is disjoint, and #$\#CC(X)<\infty,$ therefore there exists $j_{0}\in\mathbb{N}$ such that $X_{j}\cap X=\emptyset~{}\forall j>j_{0}.$ Hence $X_{j}=\widehat{X_{j}\cap X}=\emptyset~{}\forall j>j_{0}.$ This is a contradiction to the fact that #$\#CC(\widehat{X})=\infty.$ Next, we assume that #$\#CC(X)=m$ and #$\#CC(\widehat{X})=l.$ Assume that $\widehat{X}=\cup^{l}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ Then $X=\widehat{X}\cap X=\cup^{l}_{j=1}X_{j}\cap X.$ By using Result 2.11, we obtain that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ If each $X_{j}\cap X$ is connected, then $l=m.$ Otherwise $l<m.$
We write
$P(z_{1},z_{2})=\sum_{\begin{subarray}{c}0\leq i\leq m\\ 0\leq j\leq n\\ \end{subarray}}a_{ij}z_{1}^{i}z_{2}^{j}.$
Therefore, on $\Gamma_{\mathfrak{D}_{2}},$ we have
$\overline{(P\circ\Psi)(z_{1},z_{2})}$ $=\overline{P(p_{1},p_{2})}=\sum_{\begin{subarray}{c}0\leq i\leq m\\ 0\leq j\leq n\\ \end{subarray}}\overline{a}_{ij}\overline{p_{1}}^{i}\overline{{p_{2}}}^{j}=\frac{1}{p^{m}_{1}p^{n}_{2}}\sum_{\begin{subarray}{c}0\leq i\leq m\\ 0\leq j\leq n\\ \end{subarray}}\overline{a}_{ij}p_{1}^{m-i}{p_{2}}^{n-j}$
$=\frac{(K\circ\Psi)(z_{1},z_{2})}{p^{m}_{1}p^{n}_{2}}=(h\circ\Psi)(z_{1},z_{2}),$
where
$K(z_{1},z_{2})=\sum_{\begin{subarray}{c}0\leq i\leq m\\ 0\leq j\leq n\\ \end{subarray}}\overline{a}_{ij}z_{1}^{m-i}z_{2}^{n-j}\text{ and }\quad h(z_{1},z_{2})=\frac{K(z_{1},z_{2})}{z^{m}_{1}z^{n}_{2}}.$
We define
$L:=$$\{p_{1}=0,|p_{2}|\leq 1\}\cup\{p_{2}=0,|p_{1}|\leq 1\}\text{ and }$
$X:=$$\left\{(z_{1},z_{2})\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{(P\circ\Psi)(z_{1},z_{2})}=(h\circ\Psi)(z_{1},z_{2})\right\}.$
Lemma 4.2.
$X$ is totally real at $z_{0}\in X,$ if
$\triangle({z_{0}}):=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}\neq 0.$
Proof.
Take
$\rho_{1}$ $={\sf Re}\left(\overline{(P\circ\Psi)(z_{1},z_{2})}-(h\circ\Psi)(z_{1},z_{2})\right),$
$\rho_{2}$ $={\sf Im}\left(\overline{(P\circ\Psi)(z_{1},z_{2})}-(h\circ\Psi)(z_{1},z_{2})\right).$
Therefore,
$\rho_{1}(z)$ $=\frac{1}{2}\left(\overline{(P\circ\Psi)(z)}-(h\circ\Psi)(z)+(P\circ\Psi)(z)-\overline{(h\circ\Psi)(z)}\right)$
$\rho_{2}(z)$ $=\frac{1}{2i}\left(\overline{(P\circ\Psi)(z)}-(h\circ\Psi)(z)-(P\circ\Psi)(z)+\overline{(h\circ\Psi)(z)}\right).$
Since
$X=\left\{(z,w)\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\rho_{1}=0,\rho_{2}=0\right\},$
$X$ is totally real at $z_{0}\in X$ if
$\begin{vmatrix}\frac{\partial\rho_{1}}{\partial{{z_{1}}}}&\frac{\partial\rho_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial\rho_{2}}{\partial{{z_{1}}}}&\frac{\partial\rho_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}\neq 0.$
We compute:
$\frac{\partial\rho_{1}}{\partial{z_{j}}}$ $=\frac{1}{2}\left(-\frac{\partial(h\circ\Psi)}{\partial{z_{j}}}+\frac{\partial(P\circ\Psi)}{\partial{z_{j}}}\right)$
$\frac{\partial\rho_{2}}{\partial{z_{j}}}$ $=\frac{1}{2i}\left(-\frac{\partial(h\circ\Psi)}{\partial{z_{j}}}-\frac{\partial(P\circ\Psi)}{\partial{z_{j}}}\right),$
for $j=1,2.$ And
$\frac{\partial(h\circ\Psi)}{\partial{z_{j}}}$ $=\frac{\partial h(\Psi)}{\partial{z_{1}}}\frac{\partial p_{1}}{\partial{z_{j}}}+\frac{\partial h(\Psi)}{\partial{z_{2}}}\frac{\partial p_{2}}{\partial{z_{j}}}$
$\frac{\partial(P\circ\Psi)}{\partial{z_{j}}}$ $=\frac{\partial P(\Psi)}{\partial{z_{1}}}\frac{\partial p_{1}}{\partial{z_{j}}}+\frac{\partial P(\Psi)}{\partial{z_{2}}}\frac{\partial p_{2}}{\partial{z_{j}}},~{}j=1,2.$
Hence
$\begin{vmatrix}\frac{\partial\rho_{1}}{\partial{{z_{1}}}}&\frac{\partial\rho_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial\rho_{2}}{\partial{{z_{1}}}}&\frac{\partial\rho_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}=$ $\begin{vmatrix}\frac{1}{2}\left(-\frac{\partial(h\circ\Psi)}{\partial{z_{1}}}+\frac{\partial(P\circ\Psi)}{\partial{z_{1}}}\right)&\frac{1}{2}\left(-\frac{\partial(h\circ\Psi)}{\partial{z_{2}}}+\frac{\partial(P\circ\Psi)}{\partial{z_{2}}}\right)\\[6.45831pt] \frac{1}{2i}\left(-\frac{\partial(h\circ\Psi)}{\partial{z_{1}}}+\frac{\partial(P\circ\Psi)}{\partial{z_{1}}}\right)&\frac{1}{2i}\left(-\frac{\partial(h\circ\Psi)}{\partial{z_{2}}}+\frac{\partial(P\circ\Psi)}{\partial{z_{2}}}\right)\\[6.45831pt] \end{vmatrix}_{z_{0}}$
$=$ $\frac{1}{2i}\begin{vmatrix}-\frac{\partial(h\circ\Psi)}{\partial{z_{1}}}&-\frac{\partial(h\circ\Psi)}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial(P\circ\Psi)}{\partial{z_{1}}}&\frac{\partial(P\circ\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}.$
Therefore,
$\begin{vmatrix}\frac{\partial\rho_{1}}{\partial{{z_{1}}}}&\frac{\partial\rho_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial\rho_{2}}{\partial{{z_{1}}}}&\frac{\partial\rho_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}\neq 0\text{ if and only if }$ $\begin{vmatrix}\frac{\partial(P\circ\Psi)}{\partial{z_{1}}}&\frac{\partial(P\circ\Psi)}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial(h\circ\Psi)}{\partial{z_{1}}}&\frac{\partial(h\circ\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}$ $=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}\times\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z_{0}}\neq 0.$
We can write
$\triangle(z)=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}\times\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z}=\frac{1}{p^{m+1}_{1}p^{n+1}_{2}}\Pi^{l}_{j=1}q_{j}(z),$
for some $l\in\mathbb{N},$ where each $q_{j}$ is an irreducible polynomial in $\mathbb{C}^{2}.$ We define the corresponding irreducible algebraic variety $Z_{j}:=\{z\in\mathbb{C}^{2}:q_{j}(z)=0\}.$ We assume $\triangle({z})\not\equiv 0$ on $X.$ Therefore, each $q_{j}$ is a non-zero holomorphic polynomial in $\mathbb{C}^{2}.$ We set
$\Sigma:=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\triangle{(z)}=0\right\}.$ (4.1)
We denote
$Q_{j}=Z_{j}\cap\Gamma_{\mathfrak{D}_{2}}.$ (4.2)
Theorem 4.3.
We let $J=\{j\in\{1,\cdots,l\}:\emptyset\neq Q_{j}\neq\widehat{Q_{j}},\widehat{Q_{j}}\setminus L\subset X\}.$
  1. (i)
    If $J=\emptyset,$ then $\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}}).$
  2. (ii)
    If $J\neq\emptyset,$ then
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})\cup\bigg{(}\cup_{j\in J}{\sf Gr}_{\overline{P\circ\Psi}}(\widehat{Q_{j}})\bigg{)}.$
Remark 4.4.
If (i) of Theorem 4.3 is true, then by Theorem 2.24, we get that
$[z_{1},z_{2},\overline{P\circ\Psi}]_{\Gamma_{{\mathfrak{D}_{2}}}}=\mathcal{C}(\Gamma_{{\mathfrak{D}_{2}}}).$
We will prove this theorem through some lemmas.
Lemma 4.5.
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\overline{\mathfrak{D}_{2}}})}\subset{\sf Gr}_{\overline{P\circ\Psi}}(X\cap\Sigma)\cup{\sf Gr}_{\overline{P\circ\Psi}}(L\cup\Gamma_{\mathfrak{D}_{2}}),$ where $\Sigma$ as in (4.1).
Proof.
First, we show that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{\overline{P\circ\Psi}}(X)\cup{\sf Gr}_{\overline{P\circ\Psi}}(L\cup\Gamma_{\mathfrak{D}_{2}}).$
Since $\overline{\mathfrak{D}_{2}}$ is polynomially convex, $\mathfrak{D}_{2}$ has the segment property, and ${\overline{P\circ\Psi}}$ is pluriharmonic function, therefore using Lemma 2.21, we can say that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{\overline{P\circ\Psi}}(\overline{\mathfrak{D}_{2}}).$ (4.3)
We now compute:
${\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})$ =$=\left\{(z_{1},z_{2},\xi):\xi=\overline{(P\circ\Psi)(z_{1},z_{2})},(z_{1},z_{2})\in\Gamma_{\mathfrak{D}_{2}}\right\}$
=$=\left\{(z_{1},z_{2},\xi):\xi=\overline{(P\circ\Psi)(z_{1},z_{2})}=(h\circ\Psi)(z_{1},z_{2}),(z_{1},z_{2})\in\Gamma_{\mathfrak{D}_{2}}\right\}$
=$=\left\{(z_{1},z_{2},\xi):\xi=(h\circ\Psi)(z_{1},z_{2})=\frac{(K\circ\Psi)(z_{1},z_{2})}{p^{m}_{1}p^{n}_{2}},(z_{1},z_{2})\in\Gamma_{\mathfrak{D}_{2}}\right\}$
=$=\left\{(z_{1},z_{2},\xi):\xi{p^{m}_{1}p^{n}_{2}}-{(K\circ\Psi)(z_{1},z_{2})}=0,(z_{1},z_{2})\in\Gamma_{\mathfrak{D}_{2}}\right\}.$
Therefore,
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\subset\left\{(z_{1},z_{2},\xi):\xi{p^{m}_{1}p^{n}_{2}}-{(K\circ\Psi)(z_{1},z_{2})}=0,(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}\right\}.$ (4.4)
Using (4.3) and (4.4), we get that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{\overline{P\circ\Psi}}(X)\cup{\sf Gr}_{\overline{P\circ\Psi}}(L\cup\Gamma_{\mathfrak{D}_{2}}).$
Again, since by Lemma 4.2, $X\setminus\Sigma$ is totally real, using Lemma 2.23, we obtain that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{\overline{P\circ\Psi}}(X\cap\Sigma)\cup{\sf Gr}_{\overline{P\circ\Psi}}(L\cup\Gamma_{\mathfrak{D}_{2}}),$
which completes the proof. ∎
By Lemma 4.1, we have that the number of connected component of $\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}$ is less than or equal to the number of connected components of ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}}).$ Therefore, it is enough to consider the set ${\sf Gr}_{\overline{P\circ\Psi}}(X\cap\Sigma).$
Lemma 4.6.
Fix $j_{0}\in\{1,\cdots,l\}.$ If there exists $\alpha\in(Z_{j_{0}})_{reg}$ such that ${\sf Gr}_{\overline{P\circ\Psi}}(\alpha)\in\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})},$ then $P\circ\Psi=c_{j_{0}},\text{ on }Z_{j_{0}}$ for some constant $c_{j_{0}}\in\mathbb{C}.$
Proof.
We choose a neighborhood $U$ of $\alpha$ such that $U\cap Z_{j_{0}}=:D_{\alpha}$ is a smoothly bounded disc containing $\alpha.$ Let $X:=\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial U\times\mathbb{C}^{N}).$ Since $\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{\mathfrak{D}_{2}}\cup[\cup^{l}_{j=1}Z_{j}]),$ we have
$X\subset{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{\mathfrak{D}_{2}}\cup[\cup^{l}_{j=1}Z_{j}])\cap(\partial U\times\mathbb{C}^{N})\subset{\sf Gr}_{\overline{(P\circ\Psi)}}(\partial U\cap Z_{j_{0}}).$
We claim that
${\sf Gr}_{\overline{(P\circ\Psi)}}(\alpha)\in\widehat{X}\subset\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\partial U\cap V)}=\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\partial D_{\alpha})}.$
To prove this claim, let $E={\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ and $N=U\times\mathbb{B}(r)$ with $r>\sup_{U}|\overline{(P\circ\Psi)}|.$ Note that $E\cap N=\emptyset$ and $\partial N\cap\widehat{E}=[(\partial U\times\mathbb{B}(r))\cap\widehat{E}]\cup[(U\times\partial\mathbb{B}(r))\cap\widehat{E}].$ We claim that $(U\times\partial\mathbb{B}(r))\cap\widehat{E}=\emptyset:$ if possible, let $(\eta,\xi)\in(U\times\partial\mathbb{B}(r))\cap\widehat{E}.$ Then $\xi=\overline{(P\circ\Psi)}(\eta)$ and $\eta\in U,|\xi|=r.$ But on $U,$ we have $|\xi|=|\overline{(P\circ\Psi)}(\eta)|<r.$ This is a contradiction to the assumption that $(U\times\partial\mathbb{B}(r))\cap\widehat{E}\not=\emptyset.$ Then by Result 2.4, we have
${\sf Gr}_{\overline{(P\circ\Psi)}}(\alpha)\in\widehat{\widehat{E}\cap\partial N}$ $=\widehat{[\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial U\times\mathbb{B}(r))]}$
$\subset\widehat{[\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{\mathfrak{D}_{2}})}\cap(\partial U\times\mathbb{C}^{N})]}$
$=\widehat{X}\subset\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\partial U\cap V)}=\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\partial D_{\alpha})}.$
Therefore, $\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\partial D_{\alpha})}$ is not polynomially convex, hence by the Wermer’s maximality theorem we can say that $\overline{(P\circ\Psi)}$ is holomorphic in $D_{\alpha}.$ If $\alpha$ runs over regular points of $Z_{j},$ the corresponding neighborhoods $D_{\alpha}$ cover $(Z_{j_{0}})_{\text{reg}}.$
Lemma 4.7.
Fix $j_{0}\in\{1,\cdots,l\}.$ If $P\circ\Psi\equiv c_{j_{0}},\text{ on }Z_{j_{0}}$ for some constant $c_{j_{0}}\in\mathbb{C}$ but $h\circ\Psi\not\equiv\bar{c}_{j_{0}}$ on $Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\setminus L,$ then
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\cap{\sf Gr}_{\overline{P\circ\Psi}}\left(X\cap Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})\right)=\emptyset.$
Proof.
Since $P\circ\Psi-c_{j_{0}}\equiv 0,\text{ on }Z_{j_{0}}$ and $q_{j_{0}}$ is an irreducible polynomial, therefore $q_{j_{0}}$ divides $P\circ\Psi-c_{j_{0}}.$ Now
$X\cap Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}=\{z\in\overline{\mathfrak{D}}_{2}\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}:q_{j_{0}}(z)=0,(h\circ\Psi)(z)-c_{j_{0}}=0)\}$
is finite. Hence
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\cap{\sf Gr}_{\overline{P\circ\Psi}}\left(X\cap Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})\right)=\emptyset.$
Remark 4.8.
In view of Lemma 4.6 and Lemma 4.7, we obtain that if there exists a regular point $\alpha\in Z_{j}$ such that ${\sf Gr}_{\overline{(P\circ\Psi)}}(\alpha)\in\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})},$ then $P\circ\Psi=c_{j}$ and $h\circ\Psi=\overline{c}_{j}$ on $Z_{j}$ for some constant $c_{j}.$
We write
$\mathfrak{J}_{0}:=\{j\in\{1,\cdots,l\}:\emptyset\not=\overline{\mathfrak{D}}_{2}\cap Z_{j}\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})\subset X\}.$
And
$P\circ\Psi=c_{j}\text{ and }h\circ\Psi=\overline{c_{j}},~{}j\in\mathfrak{J}_{0}.$ (4.5)
From above lemmas we get that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}\subset{\sf Gr}_{\overline{P\circ\Psi}}(L\cup\Gamma_{\mathfrak{D}_{2}})\cup\bigg{(}\cup_{j\in\mathfrak{J}_{0}}{\sf Gr}_{\overline{P\circ\Psi}}(Z_{j}\cap\overline{\mathfrak{D}}_{2}\cap X)\bigg{)}.$
Lemma 4.9.
If $j\in\mathfrak{J}_{0}$ and $Q_{j}$ as in (4.2), then ${\sf Gr}_{\overline{P\circ\Psi}}(\widehat{Q}_{j})\subset\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
Proof.
Let $(a,\overline{(P\circ\Psi)}(a))\in{\sf Gr}_{\overline{P\circ\Psi}}(\widehat{Q}_{j})$ and let $H$ be any polynomial in $\mathbb{C}^{3}.$ Then
$|H(a,\overline{(P\circ\Psi)(a)})|$ $\leq\sup_{{\sf Gr}_{\overline{P\circ\Psi}}(\widehat{Q}_{j})}|H(z,\xi)|$
[…]
$=\sup_{{Q}_{j}}|H(z,\overline{(P\circ\Psi)(z)})|$
$\leq\sup_{\Gamma_{\mathfrak{D}_{2}}}|H(z,\overline{(P\circ\Psi)(z)})|.$
Hence $(a,\overline{(P\circ\Psi)(a)})\in\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}.$ Therefore, ${\sf Gr}_{\overline{P\circ\Psi}}(\widehat{Q}_{j})\subset\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
Lemma 4.10.
If $j_{0}\in\mathfrak{J}_{0}$ and $Q_{j}$ as in (4.2), then ${\sf Gr}_{\overline{P\circ\Psi}}\left(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}\cup\widehat{Q}_{j_{0}})\right)\cap\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}=\emptyset.$
Proof.
We write $(P\circ\Psi)(z)-c_{j_{0}}=f_{1}(z)q^{m_{j_{0}}}_{j_{0}}(z),$ where $m_{j_{0}}$ denotes the maximal order of an irreducible factor $q_{j_{0}}.$ We define $A=\{z\in\overline{\mathfrak{D}}_{2}:f_{1}(z)=0\}.$ Let $a_{0}\in\left(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}\cup\widehat{Q}_{j_{0}})\right).$ First, we assume that $a_{0}\in Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\setminus(L\cup A\cup\Gamma_{{\mathfrak{D}_{2}}}).$ We claim that ${\sf Gr}_{\overline{P\circ\Psi}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}.$ To prove this claim, we put
$f_{2}(z)=\frac{f_{1}(z)}{f_{1}(a_{0})}.$
Since $\widehat{Q}_{j_{0}},$$\{a_{0}\}$ are disjoint polynomially convex set, there exists a neighborhood $U$ of $\widehat{Q}_{j_{0}}$ and a neighborhood $W$ of $A$, a polynomial $p_{0}$ such that
$p_{0}(a_{0})=1,\quad|p_{0}(z)f_{2}(z)|<\frac{1}{2}~{}~{}\text{ on }U,\text{ and }$
$|p_{0}(z)f_{2}(z)|<\frac{1}{2}~{}~{}\text{ on }W.$
We set $M=\sup_{\Gamma_{{\mathfrak{D}_{2}}}}|P\circ\Psi-c_{j_{0}}|$ and $K_{1}:=\{z\in\overline{\mathfrak{D}}_{2}:(P\circ\Psi)(z)-c_{j_{0}}=0\}.$ If we define
$g_{1}(z_{1},z_{2},\xi)=1-\frac{((P\circ\Psi)(z)-c_{j_{0}})(\xi-\bar{c}_{j_{0}})}{2M^{2}},$
then $g_{1}=1$ on ${\sf Gr}_{\overline{P\circ\Psi}}(K_{1}).$ Clearly, $|g_{1}|\leq 1$ on $\Gamma_{{\mathfrak{D}_{2}}}.$ But $|g_{1}|\neq 1$ on $\Gamma_{{\mathfrak{D}_{2}}}\setminus(W\cup U).$ This implies $|g_{1}|<1$ on $\Gamma_{{\mathfrak{D}_{2}}}\setminus(W\cup U).$ Therefore, there exists a natural number $k$ such that $|f_{2}(z)p_{0}(z)g_{1}(z,\xi)^{k}|<\frac{1}{2}$ on $\Gamma_{{\mathfrak{D}_{2}}}\setminus(W\cup U).$ If we take $g(z)=f_{2}(z)p_{0}(z)g_{1}(z,\xi)^{k},$ then $g((a_{0},\overline{(P\circ\Psi)}(a_{0})))=1.$ We claim that
$|g(z_{1},z_{2},\xi)|<\frac{1}{2}\text{ on }{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}}).$
Assume that $\sup_{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}|g(z_{1},z_{2},\xi)|=|g(s,\overline{P\circ\Psi}(s))|$ for some $s=(s_{1},s_{2})\in\Gamma_{{\mathfrak{D}_{2}}}.$ Case I: $s\in\Gamma_{{\mathfrak{D}_{2}}}\setminus(U\cup W),$ then there is nothing to prove. Case II: $s\in U$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$ Case III: $s\in W$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$ Combining all the three cases, we can conclude that ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$ Next, we assume that $a_{0}\in(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}).$ Since $(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})$ is finite and $\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}$ does not containing any isolated point, therefore, in this case also ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
In view of Lemma 4.10 and Lemma 4.9, we get that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})\cup\bigg{(}\cup_{j\in J}{\sf Gr}_{\overline{P\circ\Psi}}(\widehat{Q_{j}})\bigg{)}.$
5. Examples
Example 5.1.
Let $p_{1}(z_{1},z_{2})=az_{1}+bz^{l}_{2},~{}~{}p_{2}(z_{1},z_{2})=cz_{1}+dz^{l}_{2}$ $(l\in\mathbb{N},ad-bc\neq 0,|a|\neq|c|)$ and $\overline{\mathfrak{D}}_{2}:=\{|p_{1}(z)|\leq 1,|p_{2}(z)|\leq 1\}.$ Let $f_{j}\in\mathcal{C}(\Gamma_{\mathfrak{D}_{2}})$ for $j=1,\cdots,N,N\geq 1,$ and assume that each $f_{j}$ extends to a pluriharmonic function on $\mathfrak{D}_{2}.$ Then either
$[z_{1},z_{2},f_{1},\cdots,f_{N};\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}),$
or there exists a non-trivial algebraic variety $V\subset\mathbb{C}^{2}$ with $V\cap\partial\mathfrak{D}_{2}\subset\Gamma_{\mathfrak{D}_{2}},$ and the pluriharmonic extensions of the $f_{j}$ are holomorphic on $V.$ Explanation: We claim that
  1. (i)
    $\Psi=(p_{1},p_{2}):\mathbb{C}^{2}\to\mathbb{C}^{2}$ is a proper map;
  2. (ii)
    $\mathfrak{D}_{2}$ is complex non-degenerate;
  3. (iii)
    $\overline{\mathfrak{D}}_{2}$ is simply connected;
  4. (iv)
    $\{p_{1}(z)=e_{1}\},$$\{p_{2}(z)=e_{2}\}$ is isomorphic to $\mathbb{C}$ for any $e_{1},e_{2}\in\mathbb{C}.$
Proof of (i): We consider the map $g(z_{1},z_{2})=(z_{1},z^{l}_{2})$ and $\phi(z_{1},z_{2})=(az_{1}+bz_{2},cz_{1}+dz_{2}),$ then $\Psi=\phi\circ g.$ Since $g$ is proper map and $\phi$ is an automorphism of $\mathbb{C}^{2},$ $\Psi$ is proper.
Proof of (ii): We compute
$dp_{1}\wedge dp_{2}=l(ad-bc)z^{l-1}_{2}dz_{1}\wedge dz_{2}.$
[…]
$H(t,x)=\left(t^{l}\gamma_{1}(x),t\gamma_{2}(x)\right).$
Clearly, $H(t,x)\in\overline{\mathfrak{D}}_{2}~{}~{}\forall t,x\in[0,1],~{}H(0,x)\equiv\gamma_{0}$ and $H(1,x)=\left(\gamma_{1}(x),\gamma_{2}(x)\right)=\gamma(x).$ Since $\gamma$ is continuous, $H$ is continuous. Hence $\gamma$ is homotopic to $\gamma_{0}.$ Since $\gamma$ is arbitrary, $\overline{\mathfrak{D}}_{2}$ is simply connected.
Proof of (iv): Without loss of generality, assume that $|a|\neq 0\neq|c|.$ Then $p_{1}(z)=e_{1}$ and $p_{2}(z)=e_{2}$ are simply connected for any $e_{1},e_{2}\in\mathbb{C}$. To prove this we use the following change of coordinate system $(z^{*}_{1},z^{*}_{2})=(az_{1}+bz^{l}_{2}-e_{1},z_{2})$ for the first one and $(z^{*}_{1},z^{*}_{2})=(cz_{1}+dz^{l}_{2}-e_{2},z_{2})$ for the second one. Then $p_{1}(z)=e_{1}$ and $p_{2}(z)=e_{2}$ becomes $\{(z^{*}_{1},z^{*}_{2})\in\mathbb{C}^{2}:z^{*}_{1}=0\}$ and $\{(z^{*}_{1},z^{*}_{2})\in\mathbb{C}^{2}:z^{*}_{1}=0\}$ respectively. Now by using Theorem 1.2, we can conclude that either
$[z_{1},z_{2},f_{1},\cdots,f_{N};\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}),$
or there exists a non-trivial algebraic variety $V\subset\mathbb{C}^{2}$ with $V\cap\partial\mathfrak{D}_{2}\subset\Gamma_{\mathfrak{D}_{2}},$ and the pluriharmonic extensions of the $f_{j}$ are holomorphic on $V.$
Example 5.2.
Let $p_{1}(z_{1},z_{2})=z_{1}+z_{2},~{}~{}p_{2}(z_{1},z_{2})=z_{1}z_{2},~{}~{}P(z_{1},z_{2})=z_{1}+z_{2}$ and $\Psi(z_{1},z_{2})=(p_{1}(z_{1},z_{2}),p_{2}(z_{1},z_{2})).$ Then the graph of $\overline{P\circ\Psi}$ over $\Gamma_{{\mathfrak{D}_{2}}}$ is not polynomially convex. Hence polynomial approximation does not holds.
Explanation: Following the notation in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{1}{z_{2}}.$ Therefore,
$X$ =$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{(P\circ\Psi)(z)}=(h\circ\Psi)(z)\right\}$
=$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{z_{1}+z_{2}+z_{1}z_{2}}=\frac{1}{z_{1}+z_{2}}+\frac{1}{z_{1}z_{2}}\right\}.$
and
$\triangle({z})$ $=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z}=\begin{vmatrix}1&1\\ \frac{-1}{p^{2}_{1}(z)}&\frac{-1}{p^{2}_{2}(z)}\end{vmatrix}\begin{vmatrix}1&1\\ z_{2}&z_{1}\end{vmatrix}$
$=\frac{1}{p^{2}_{1}p^{2}_{2}}(p_{1}+p_{2})(p_{1}-p_{2})(z_{2}-z_{1}).$
We define $q_{1}:=p_{1}-p_{2}=z_{1}+z_{2}-z_{1}z_{2},~{}~{}q_{2}:=p_{1}+p_{2}=z_{1}+z_{2}+z_{1}z_{2},~{}~{}q_{3}=z_{2}-z_{1}$ and $Z_{j}=\{q_{j}=0\}.$ Therefore, Clearly, $Z_{j}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ only for ${j}=2.$ It is easy to see that $\widehat{Q}_{2}=\widehat{Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+z_{2}+z_{1}z_{2}=0\}.$ Indeed, we write $\mathcal{K}=\{(z_{1},z_{2})\in\mathbb{C}^{2}:z_{1}+z_{2}=0,|z_{1}|=1,|z_{2}|=1\}.$ Clearly, $\Psi^{-1}(\mathcal{K})=Q_{2}=Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ We now claim that
$\Psi^{-1}(\widehat{\mathcal{K}})=\widehat{Q_{2}}.$
Note that
$\Psi(Q_{2})\subset\Psi(\widehat{Q_{2}})\subset\Psi(Z_{2}\cap\overline{\mathfrak{D}}_{2})=\widehat{\mathcal{K}}.$
This implies
$\widehat{\mathcal{K}}=\widehat{\Psi(Q_{2})}\subset\widehat{\Psi(\widehat{Q_{2}})}\subset\Psi(Z_{2}\cap\overline{\mathfrak{D}}_{2})=\widehat{\mathcal{K}}.$
Hence $\widehat{\mathcal{K}}=\widehat{\Psi(Q_{2})}=\widehat{\Psi(\widehat{Q_{2}})}.$ Since $\Psi^{-1}(\Psi(\widehat{Q_{2}}))=\widehat{Q_{2}}$ is polynomially convex and $\Psi$ is proper map, therefore $\Psi(\widehat{Q_{2}})$ is also polynomially convex. Thus $\widehat{\Psi(Q_{2})}=\widehat{\Psi(\widehat{Q_{2}})}=\Psi(\widehat{Q_{2}}).$ This implies $\widehat{Q_{2}}=\Psi^{-1}(\widehat{\mathcal{K}}),$ i.e. $\widehat{Q_{2}}=\overline{\mathfrak{D}}_{2}\cap Z_{2}.$ Therefore, by Theorem 4.3, we get that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})\cup\{(z_{1},z_{2},0):(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2},z_{1}+z_{2}+z_{1}z_{2}=0\}.$
Example 5.3.
$p_{1}(z_{1},z_{2})=z_{1}+z_{2},~{}~{}p_{2}(z_{1},z_{2})=z_{1}z_{2},~{}~{}P(z_{1},z_{2})=z_{1}+2z_{2}$ and $\Psi(z_{1},z_{2})=(p_{1}(z_{1},z_{2}),p_{2}(z_{1},z_{2})).$ Then the graph of $\overline{P\circ\Psi}$ over $\Gamma_{{\mathfrak{D}_{2}}}$ is polynomially convex. Hence polynomial approximation holds.
Explanation: Following the notation in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{2}{z_{2}}.$ Therefore,
$X$ =$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\overline{\mathfrak{D}_{2}}}):\overline{(P\circ\Psi)(z)}=(h\circ\Psi)(z)\right\}$
=$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{z_{1}+z_{2}+2z_{1}z_{2}}=\frac{1}{z_{1}+z_{2}}+\frac{2}{z_{1}z_{2}}\right\},$
and
$\triangle({z})$ $=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z}=\begin{vmatrix}1&2\\ \frac{-1}{p^{2}_{1}(z)}&\frac{-2}{p^{2}_{2}(z)}\end{vmatrix}\begin{vmatrix}1&1\\ z_{2}&z_{1}\end{vmatrix}$
$=\frac{2}{p^{2}_{1}p^{2}_{2}}(p_{1}+p_{2})(p_{1}-p_{2})(z_{2}-z_{1}).$
We define $q_{1}:=p_{1}-p_{2}=z_{1}+z_{2}-z_{1}z_{2},~{}~{}q_{2}:=p_{1}+p_{2}=z_{1}+z_{2}+z_{1}z_{2},~{}~{}q_{3}=z_{2}-z_{1}$ and $Z_{j}=\{q_{j}=0\}.$ It is easy see that $Z_{j}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\nsubseteq X~{}\forall{j}\in\{1,2,3\}.$ Hence,
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\overline{\mathfrak{D}_{2}}}).$
Therefore, by using Theorem 2.24, we get that $[z_{1},z_{2},\overline{(P\circ\Psi)}(z);\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$
Example 5.4.
Let $p_{1}(z_{1},z_{2})=2z_{1}+z^{2}_{2},~{}~{}p_{2}(z_{1},z_{2})=z_{1}-z^{2}_{2},~{}~{}P(z_{1},z_{2})=z_{1}-z_{2}$ and $\Psi(z_{1},z_{2})=(p_{1}(z_{1},z_{2}),p_{2}(z_{1},z_{2})).$ Then the graph of $\overline{P\circ\Psi}$ over $\Gamma_{{\mathfrak{D}_{2}}}$ is not polynomially convex. Hence polynomial approximation does not holds.
Explanation: Following the notation in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}-\frac{1}{z_{2}}.$ Therefore,
$X$ =$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{(P\circ\Psi)(z)}=(h\circ\Psi)(z)\right\}$
=$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{z_{1}+2z^{2}_{2}}=\frac{1}{2z_{1}+z^{2}_{2}}+\frac{1}{z_{1}-z^{2}_{2}}\right\}.$
and
$\triangle({z})$ $=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z}=\begin{vmatrix}1&-1\\ \frac{-1}{p^{2}_{1}(z)}&\frac{1}{p^{2}_{2}(z)}\end{vmatrix}\begin{vmatrix}2&2z_{2}\\ 1&-2z_{2}\end{vmatrix}$
$=\frac{1}{p^{2}_{1}p^{2}_{2}}(p_{1}+p_{2})(p_{1}-p_{2})(-6z_{2}).$
We define $q_{1}:=p_{1}-p_{2}=z_{1}+2z^{2}_{2},~{}~{}q_{2}:=p_{1}+p_{2}=3z_{1},~{}~{}q_{3}=-6z_{2}$ and $Z_{j}=\{q_{j}=0\}.$ Clearly, $Q_{3}=Z_{3}\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset$ and it is easy to show that if $z\in Z_{2}\cap X,$ then $z\in\Gamma_{{\mathfrak{D}_{2}}},$ this is not possible because $X\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset.$ Now its remaining to verify for the set $Q_{1}=Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ It is easy to see that $Z_{1}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ and $\widehat{Q}_{1}=\widehat{Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+2z^{2}_{2}=0\}.$ Therefore, by Theorem 4.3, we get that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})\cup\{(z_{1},z_{2},0):(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2},z_{1}+2z^{2}_{2}=0\}.$
Example 5.5.
Let $p_{1}(z_{1},z_{2})=2z_{1}+z^{2}_{2},~{}~{}p_{2}(z_{1},z_{2})=z_{1}-z^{2}_{2},~{}~{}P(z_{1},z_{2})=z_{1}+2z_{2}$ and $\Psi(z_{1},z_{2})=(p_{1}(z_{1},z_{2}),p_{2}(z_{1},z_{2})).$ Then the graph of $\overline{P\circ\Psi}$ over $\Gamma_{{\mathfrak{D}_{2}}}$ is polynomially convex.
Explanation: Following the notation in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{2}{z_{2}}.$ Therefore,
$X$ =$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{(P\circ\Psi)(z)}=(h\circ\Psi)(z)\right\}$
=$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{4z_{1}-z^{2}_{2}}=\frac{1}{2z_{1}+z^{2}_{2}}+\frac{2}{z_{1}-z^{2}_{2}}\right\}.$
and
$\triangle({z})$ $=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z}=\begin{vmatrix}1&2\\ \frac{-1}{p^{2}_{1}(z)}&\frac{-2}{p^{2}_{2}(z)}\end{vmatrix}\begin{vmatrix}2&2z_{2}\\ 1&-2z_{2}\end{vmatrix}$
$=\frac{1}{p^{2}_{1}p^{2}_{2}}(p_{1}+p_{2})(p_{1}-p_{2})(6z_{2}).$
We define $q_{1}:=p_{1}-p_{2}=z_{1}+2z^{2}_{2},~{}~{}q_{2}:=p_{1}+p_{2}=3z_{1},~{}~{}q_{3}=z_{2}$ and $Z_{j}=\{q_{j}=0\}.$ Clearly, $Q_{3}=Z_{3}\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset$ and it is easy to show that if $z\in Z_{j}\cap X,$ then $z\in\Gamma_{{\mathfrak{D}_{2}}},$ for $j=1,2.$ This is not possible because $X\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset.$ Therefore, by Theorem 4.3, we get that
$\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}}).$
Therefore, by using Theorem 2.24, we get $[z_{1},z_{2},\overline{(P\circ\Psi)}(z);\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$ Acknowledgments. The first named author was partially supported by a MATRICS Research Grant (MTR/2017/000974) of SERB, Dept. of Science and Technology, Govt. of India, for the beginning of this work and is supported partially by a Core Research Grant (CRG/2022/003560) of SERB, Dept. of Science and Technology, Govt. of India, for the later part of the work. The second named author acknowledges partial support from an INSPIRE Fellowship (IF 160487) awarded by the DST, Govt. of India, for the early version of this work, and an IoE-IISc Postdoctoral Fellowship (R(HR)(IOE-PDF)(MA)(GMM)-114) for the latest version of this work. The second author would also like to thank Amar Deep Sarkar for the fruitful discussions.

### v2

@@
 Abstract.
-In this paper we extend the dichotomy given by Samuelsson and Wold [28] that can be thought of as an analogue of the Wermer maximality theorem in $\mathbb{C}^{2}$ for certain polynomial polyhedra.
+In this paper we extend the dichotomy given by Samuelsson and Wold that can be thought of as an analogue of the Wermer maximality theorem in $\mathbb{C}^{2}$ for certain polynomial polyhedra.
 We consider complex non-degenerate simply connected polynomial polyhedra of the form $\Omega:=\{z\in\mathbb{C}^{2}:|p_{1}(z)|<1,|p_{2}(z)|<1\}$ such that $\overline{\Omega}$ is compact.
@@
 Here, we treat the case when $f$ is the restriction of an anti-holomorphic polynomial in $\mathbb{C}^{2}$ to the distinguished boundary $\Gamma_{{\mathfrak{D}_{2}}}$ of a complex non-degenerate polynomial polyhedron $\mathfrak{D}_{2}$ in $\mathbb{C}^{2}.$ We present explicit descriptions of the hull of the graph of $f$ on $\Gamma_{{\mathfrak{D}_{2}}}.$ In this setting, we show that the set $\widehat{{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})}\setminus{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})$ is contained in an algebraic variety.
-We now look at the graph over certain bounded domains whose closure is polynomially convex.
-Specifically, we look at domains with polynomially convex closure and having segment property.
+We now look at the graph over certain bounded domains whose closure is polynomially convex. Specifically, we look at domains with polynomially convex closure and having segment property.
 Definition 1.5.
@@
 For more details on the segment property we refer the reader to [7].
-We now present a couple of observations that will generalize results from ([28, Theorem 1.3]) and [17, Theorem 1.1] respectively:
+We now present a couple of observations that will generalize results from ([30, Theorem 1.3]) and [17, Theorem 1.1] respectively:
 Theorem 1.6.
@@
     $[z,h;\overline{\Omega}]=\left\{f\in\mathcal{C}(\overline{\Omega}):f|_{\partial\Omega}\in[z,h;\partial\Omega]\right\}.$
-We observe that the $\mathcal{C}^{1}$-smoothness of $\Omega$ is required to prove ([28, Theorem 1.3]) and ([17, Theorem 1.1]) only to show that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex.
+We observe that, in ([17, Theorem 1.1]) and ([30, Theorem 1.3]), the $\mathcal{C}^{1}$-smoothness of $\Omega$ is required to prove and only to show that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex.
 By Lemma 2.21, we can say that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex for any bounded domain $\Omega$ with polynomially convex closure and having segment property except possibly at a countable set of boundary points.
-The rest follows from [28, Theorem 1.3] and [17, Theorem 1.1], respectively.
+The rest follows from [30, Theorem 1.3] and [17, Theorem 1.1], respectively.
+Although the proof of Theorem 1.6 and Theorem 1.7 are not very different from that of [17] except for the polynomial convexity, but there are huge implication of Theorem 1.6 and Theorem 1.7 as we can deal with a huge class of nonsmooth domains here.
+A domain having continuous boundary satisfies the segment property.
+Although whether the closure of a bounded domain is polynomially convex is a very difficult problem, the class of bounded convex domains satisfies the segment property and their closure is also polynomially convex.
+Hence, we obtain the following corollary:
+Corollary 1.8.
+Let $\Omega$ be a bounded convex domain in $\mathbb{C}^{n}$.
+Let $h_{j}$, $j=1,\dots,N$, be pluriharmonic in $\Omega$ and continuous on $\overline{\Omega}$.
+Assume further that there does not exist a nontrivial analytic disc $\triangle\subset\Omega$ on which all the $h_{j}$’s are holomorphic.
+Then
+  1. (1)
+    $G_{h}(\partial\Omega)$ is polynomially convex.
+  2. (2)
+    $\widehat{G_{h}(\partial\Omega)}\setminus G_{h}(\partial\Omega)$ contains no analytic disc.
+  3. (3)
+    $[z,h;\overline{\Omega}]=\left\{f\in\mathcal{C}(\overline{\Omega}):f|_{\partial\Omega}\in[z,h;\partial\Omega]\right\}.$
+We now provide some remarks: the first two of them concerns about the comparison between Theorem 1.2 with known results on polynomial polyhedra and other domains in $\mathbb{C}^{2}$; and the other concerns about the possibility of extending Theorem 1.2 to domains in $\mathbb{C}^{n}$, $n\geq 3$.
+These comments might be interesting in the context of further research.
+Remark 1.9.
+The conditions on polynomial polyhedra in Theorem 1.2 are quite natural.
+On simply connected domain any complex valued pluriharmonic function can be seen as a sum of a holomorphic function and the real part of another holomorphic function.
+This is used crucially in our proofs.
+Complex non-degeneracy of the polyhedra gives us the segment property which is one other essential ingredient in our proof.
+A natural class of polynomial polyhedra is special polynomial polyhedra in the sense of [21].
+The class of polyhedra that satisfies conditions of Theorem 1.2 has nontrivial intersection with the class of special polynomial polyhedra.
+Any polynomially convex set satisfying some conditions (see Corollary 2 and Theorem 3 of [21]) can be approximated by special polynomial polyhedra.
+It is not clear whether Theorem 1.2 holds for special polynomial polyhedra defined in [21].
+Also, there are some approximation results of Mergelyan type on certain polynomial polyhedra (see [23], [24]).
+Remark 1.10.
+When one considers a general domain in $\mathbb{C}^{2}$, it is already difficult to find the Shilov boundary.
+Therefore, one assumes the boundary approximation as in Theorem 1.6.
+If Shilov boundary can be found precisely and it is two real dimension and the other parts of the boundary are foliated by analytic discs, then the question of getting a dichotomy as in Theorem 1.2 make sense and is also interesting.
+For certain proper holomorphic images of the bidisc such dichotomy were shown in [11].
+Remark 1.11.
+In complex dimension three or higher, even for the polydisc, things are complicated.
+Main obstruction comes in giving a stratification of the graph over the boundary.
+This is the reason for assuming boundary approximation while considering the general domains in Theorem 1.6.
+Here, unlike the case in Theorem 1.2, one need to deal with $n-1$ dimensional analytic varieties; some part of our techniques fail even in $\mathbb{C}^{3}$.
+The case of polynomial polyhedra could be thought of as the first step towards it.
 We now give a brief outline for our proofs of Theorem 1.2.
@@
 2. Technical Results
-We make use of the following result due to Samuelsson-Wold [28, Corollary 4.2].
+We make use of the following result due to Samuelsson-Wold [30, Corollary 4.2].
 Result 2.1 (Samuelsson-Wold).
@@
 Result 2.5.
-([28, Propostion 4.7]) Let $K\subset\mathbb{C}^{N}$ be compact, let $F:\mathbb{C}^{n}\to\mathbb{C}^{m}$ be the uniform limit on $K$ of entire functions, and let $Y=F(K);$ note that $F$ extends to $\widehat{K}.$ If $y\in Y$ is a peak point for the algebra $\mathscr{P}(Y),$ then $\widehat{K}\cap F^{-1}(y)=\widehat{(K\cap F^{-1}\{y\}}).$
+([30, Propostion 4.7]) Let $K\subset\mathbb{C}^{N}$ be compact, let $F:\mathbb{C}^{n}\to\mathbb{C}^{m}$ be the uniform limit on $K$ of entire functions, and let $Y=F(K);$ note that $F$ extends to $\widehat{K}.$ If $y\in Y$ is a peak point for the algebra $\mathscr{P}(Y),$ then $\widehat{K}\cap F^{-1}(y)=\widehat{(K\cap F^{-1}\{y\}}).$
 Let ${\sf psh}(\Omega)$ denote the class of plurisubharmonic functions in $\Omega$ and ${\sf psh}(\overline{\Omega})$ denotes the collection of function $\phi$ such that there exists an open set $U\ \supset\Omega$ with $\phi\in{\sf psh}(U)$.
-Sibony [29, Theorem 2.2] showed that if $\Omega$ is a smooth pseudoconvex domain, then every functions in ${\sf psh}(\Omega)\cap C(\overline{\Omega})$ can be approximated uniformly with functions in ${\sf psh}(\overline{\Omega})\cap C^{\infty}(\overline{\Omega})$ and we say $\Omega$ has the PSH-Mergelyan property.
+Sibony [31, Theorem 2.2] showed that if $\Omega$ is a smooth pseudoconvex domain, then every functions in ${\sf psh}(\Omega)\cap C(\overline{\Omega})$ can be approximated uniformly with functions in ${\sf psh}(\overline{\Omega})\cap C^{\infty}(\overline{\Omega})$ and we say $\Omega$ has the PSH-Mergelyan property.
 Fornæss and Wiegerinck [9, Theorem 1] proved that this is also true for any bounded $\mathcal{C}^{1}$-smooth domain.
@@
 if a complex algebraic curve in $\mathbb{C}^{2}$ is isomorphic to $\mathbb{C}$ then in some algebraic coordinates $(z^{*}_{1},z^{*}_{2})$ on $\mathbb{C}^{2}$ its equation is $z^{*}_{1}=0$ (see Abhyankar-Moh [1], and Suzuki [33]).
-The following result [27, Theorem 5.1] gives a criterion for a homogeneous polynomial map from $\mathbb{C}^{n}$ to $\mathbb{C}^{n}$ to be proper holomorphic.
+The following result [29, Theorem 5.1] gives a criterion for a homogeneous polynomial map from $\mathbb{C}^{n}$ to $\mathbb{C}^{n}$ to be proper holomorphic.
 Result 2.9 (Rudin).
@@
 Let $\Omega\subset\mathbb{C}^{n}$ be a domain, let $h_{j}\in PH(\Omega)$ for $j=1,\cdots,N,$ and $Z\subset\Omega$ be an irreducible analytic set of dimension $d\geq 1.$ Let $1\leq m\leq d,$ fix $(i_{1},\cdots,i_{m}),$ and define
-[…]
+$Z^{\prime}:=\{z\in Z_{\text{reg}}:i^{*}_{Z_{\text{reg}}}(\overline{\partial}h_{j_{1}}\wedge\cdots\wedge\overline{\partial}h_{j_{n}})(z)=0\}.$
 Then $\widetilde{Z}:=Z^{{}^{\prime}}\cup Z_{sing}$ is an analytic subset of $\Omega.$
@@
 If $Z$ is an analytic subvariety in $M$ then $f(Z)$ is also an analytic subvariety in $N.$ Moreover, if $Z$ is irreducible then $f(Z)$ is also irreducible subvariety of $N.$
-The following is known as Tornehave’s result ([30, Corollary 3.8.11]).
+The following is known as Tornehave’s result ([32, Corollary 3.8.11]).
 Result 2.13 (Tornehave).
@@
 We denote $\Sigma_{j}:=\{z\in\mathbb{C}^{n}:\psi_{j}(z)=0\}$ and $\Sigma_{J_{k}}:=\Sigma_{j_{1}}\cap\cdots\cap\Sigma_{j_{k}}$, where $J_{k}=({j_{1},\cdots,j_{k}})$ with $1\leq j_{1}<j_{2}<\ldots<j_{k}\leq N$.
-The topological boundary of $\Omega$ is denoted by $\partial\Omega$ and clearly $\partial\Omega\subset\cup_{j}\Sigma_{j}.$ The result presented below describes a class of domains where the segment property holds.
-A special case for Weil polyhedra was proved in [22, Proposition 2.1].
+The topological boundary of $\Omega$ is denoted by $\partial\Omega$ and clearly $\partial\Omega\subset\cup_{j}\Sigma_{j}.$
+The result presented below describes a class of domains where the segment property holds. A special case for Weil polyhedra was proved in [23, Proposition 2.1].
 Proposition 2.17.
@@
 Note that $\overline{\mathfrak{D}}_{2}\subset\{z\in\mathbb{C}^{2}:z_{1}\neq 0\}.$ Let $\gamma:=\{(e^{i\theta},e^{-i\theta}):0\leq\theta\leq 2\pi\}$ be a loop in $\overline{\mathfrak{D}}_{2}$ and take the projection map $\Pi:\{z\in\mathbb{C}^{2}:z_{1}\neq 0\}\to\mathbb{C}^{*}$ by $(z_{1},z_{2})\to z_{1}.$ Clearly, $\Pi(\gamma)=\{e^{i\theta}:0\leq\theta\leq 2\pi\}$ is not null homotopic in $\mathbb{C}^{*}.$ Hence, $\gamma$ is not null homotopic in $\overline{\mathfrak{D}}_{2}.$ Therefore, $\overline{\mathfrak{D}}_{2}$ is not simply connected.
-The proof of the following result for a $C^{1}$-smooth domain is given in Samuelsson and Wold [28, Lemma 4.5].
+The proof of the following result for a $C^{1}$-smooth domain is given in Samuelsson and Wold [30, Lemma 4.5].
 Lemma 2.21.
@@
 Since $S\subset V_{Sing}$ and the set of regular points of an irreducible variety is connected, therefore we can say that $V\setminus S$ is connected.
-But right hand side of (2.13) is disconnected, and this is a contradiction to the assumption that $V\nsubseteq Z_{1},V\nsubseteq Z_{2}$ but $V\subset Z_{1}\cup Z_{2}.$ Without loss of generality, we assume that $V\subset Z_{1}.$ Next, we claim that $V=\mathfrak{D}_{n}\cap Z_{1}.$ Assume that $V\subset Z_{1}\cap\mathfrak{D}_{n}.$ Since $Z_{1}$ is an irreducible variety and if $V$ is proper subset of $Z_{1},$ then $\dim_{\mathbb{C}}V<\dim_{\mathbb{C}}Z_{1}$ (see Izzo [3], Federer [8, Theorem 3.4.8, assertion (15)]).
+But right hand side of (2.10) is disconnected, and this is a contradiction to the assumption that $V\nsubseteq Z_{1},V\nsubseteq Z_{2}$ but $V\subset Z_{1}\cup Z_{2}.$ Without loss of generality, we assume that $V\subset Z_{1}.$
+Next, we claim that $V=\mathfrak{D}_{n}\cap Z_{1}.$ Assume that $V\subset Z_{1}\cap\mathfrak{D}_{n}.$ Since $Z_{1}$ is an irreducible variety and if $V$ is proper subset of $Z_{1},$ then $\dim_{\mathbb{C}}V<\dim_{\mathbb{C}}Z_{1}$ (see Izzo [3], Federer [8, Theorem 3.4.8, assertion (15)]).
 This is not possible, as otherwise, $V$ will be discrete.
@@
 =$=\{(z_{1},\cdots,z_{n},w_{1},\cdots,w_{n}):w_{j}p_{j}(z)=1,|p_{j}(z)|=1,j=1,\cdots,n\}\subset\mathbb{C}^{2n}.$
-We claim that $K$ is polynomially convex.
-To prove this claim, we assume that $(a,b)\in\mathbb{C}^{2n}\setminus K,$ where $a=(a_{1},\cdots,a_{n})$ and $b=(b_{1},\cdots,b_{n}).$ Case I: $b_{l}p_{l}(a)\not=1,$ $|p_{l}(a)|=1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=w_{l}p_{l}(z)-1.$ Then
+We claim that $K$ is polynomially convex. To prove this claim, we assume that $(a,b)\in\mathbb{C}^{2n}\setminus K,$ where $a=(a_{1},\cdots,a_{n})$ and $b=(b_{1},\cdots,b_{n}).$
+Case I: $b_{l}p_{l}(a)\not=1,$ $|p_{l}(a)|=1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=w_{l}p_{l}(z)-1.$ Then
 $|\psi(a,b)|>0=\sup_{K}|\psi|.$
-Hence, in this case, $(a,b)\notin\widehat{K}.$ Case II: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|<1$ for some $l\in\{1,\cdots,n\}.$ This implies $|b_{l}|>1.$ Consider the polynomial $\psi(z,w)=w_{l}.$ Then
+Hence, in this case, $(a,b)\notin\widehat{K}.$
+Case II: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|<1$ for some $l\in\{1,\cdots,n\}.$ This implies $|b_{l}|>1.$ Consider the polynomial $\psi(z,w)=w_{l}.$ Then
 $|b_{l}|=|\psi(a,b)|>1=\sup_{K}|\psi|.$
-Hence, in this case, $(a,b)\notin\widehat{K}.$ Case III: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|>1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=p_{l}(z).$ Then
+Hence, in this case, $(a,b)\notin\widehat{K}.$
+Case III: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|>1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=p_{l}(z).$ Then
 $|p_{l}(a)|=|\psi(a,b)|>1=\sup_{K}|\psi|.$
-Hence, in this case, $(a,b)\notin\widehat{K}.$ Combining all three cases, we conclude that $K$ is polynomially convex.
+Hence, in this case, $(a,b)\notin\widehat{K}.$
+Combining all three cases, we conclude that $K$ is polynomially convex.
 Since $dp_{1}\wedge\cdots\wedge dp_{n}(z)\neq 0\text{ on }\Gamma,$ $\Gamma$ is totally real, and being graph over a totally real manifold, $K$ is also totally real.
@@
 3. Proof of Theorem 1.2
-Let $h_{j}$ denote the pluriharmonic extension of $f_{j}$ to $\mathfrak{D}_{2},$ and write $h=(h_{1},\cdots,h_{N}):\overline{\mathfrak{D}}_{2}\to\mathbb{C}^{N}.$ The following two cases may occur: Case I: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is polynomially convex.
-In view of Theorem 2.24, we have that
+Let $h_{j}$ denote the pluriharmonic extension of $f_{j}$ to $\mathfrak{D}_{2},$ and write $h=(h_{1},\cdots,h_{N}):\overline{\mathfrak{D}}_{2}\to\mathbb{C}^{N}.$ The following two cases may occur:
+Case I: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is polynomially convex. In view of Theorem 2.24, we have that
 $[z_{1},z_{2},f_{1},\cdots,f_{N};\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$
@@
 Case II: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex.
-By Lemma 2.22, it is enough to find an irreducible analytic variety $V\subset\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ Since ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, there exists $\alpha\in\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Now, we discuss different situations through some lemmas.
+By Lemma 2.22, it is enough to find an irreducible analytic variety $V\subset\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$
+Since ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, there exists $\alpha\in\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Now, we discuss different situations through some lemmas.
 Lemma 3.1.
@@
 $|\psi(z_{1},z_{2})|$ $=\frac{1}{2}\left|p_{1}(\alpha)+p_{1}(z_{1},z_{2})\right|.$
-Let $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}).$ We will find $\max|\omega|$ on $\Gamma_{\mathfrak{D}_{2}},$ i.e., we need to find $\max|\omega|$ on $|\omega-p_{1}(\alpha)|=|p_{1}(z)|=1.$ Let $\omega=x+iy~{}~{}\text{ and }p_{1}(\alpha)=\lambda_{1}+i\lambda_{2}.$ The equation of the straight line through the origin and through the centre of the circle $|\omega-p_{1}(\alpha)|=1$ is given by $y=\frac{\lambda_{2}}{\lambda_{1}}x.$ Solving $y=\frac{\lambda_{2}}{\lambda_{1}}x$ and $(x-\lambda_{1})^{2}+(y-\lambda_{2})^{2}=1,$ we get that $(x,y)=(0,0),~{}~{}\text{ or }(x,y)=(2\lambda_{1},2\lambda_{2}).$ Therefore, $\max|\omega|=2$ and maximum attains at $\omega=2\lambda_{1}+i2\lambda_{2}=2p_{1}(\alpha).$ Then $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}),$ implies $p_{1}(z)=p_{1}(\alpha).$ Therefore, $\max|\psi|=1$ on $\Gamma_{\mathfrak{D}_{2}}$ and since $|\omega|<1$ if $\omega\not=2p_{1}(\alpha),$ $|\psi|<1\text{ on }\Gamma_{\mathfrak{D}_{2}}\setminus\partial\triangle_{\alpha}.$ Finally we define the polynomial in $\mathbb{C}^{2+N}$
+Let $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}).$ We will find $\max|\omega|$ on $\Gamma_{\mathfrak{D}_{2}},$ i.e., we need to find $\max|\omega|$ on $|\omega-p_{1}(\alpha)|=|p_{1}(z)|=1.$ Let $\omega=x+iy~{}~{}\text{ and }p_{1}(\alpha)=\lambda_{1}+i\lambda_{2}.$ The equation of the straight line through the origin and through the centre of the circle $|\omega-p_{1}(\alpha)|=1$ is given by $y=\frac{\lambda_{2}}{\lambda_{1}}x.$ Solving $y=\frac{\lambda_{2}}{\lambda_{1}}x$ and $(x-\lambda_{1})^{2}+(y-\lambda_{2})^{2}=1,$ we get that $(x,y)=(0,0),~{}~{}\text{ or }(x,y)=(2\lambda_{1},2\lambda_{2}).$ Therefore, $\max|\omega|=2$ and maximum attains at $\omega=2\lambda_{1}+i2\lambda_{2}=2p_{1}(\alpha).$ Then $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}),$ implies $p_{1}(z)=p_{1}(\alpha).$ Therefore, $\max|\psi|=1$ on $\Gamma_{\mathfrak{D}_{2}}$ and since $|\omega|<1$ if $\omega\not=2p_{1}(\alpha),$ $|\psi|<1\text{ on }\Gamma_{\mathfrak{D}_{2}}\setminus\partial\triangle_{\alpha}.$
+Finally we define the polynomial in $\mathbb{C}^{2+N}$
 $\chi(z,w):=\psi^{m}(z)\times\rho(z,w).$
@@
 Lemma 3.2.
-[…]
+Assume that $\widetilde{V}\not=\mathfrak{D}_{2}.$ Let $V$ be an irreducible component of $\widetilde{V}$ and let $\alpha\in V\cap(\widetilde{V})_{\text{reg}}$ with ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Then ${\sf Gr}_{h}(V)\subset\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}$ and all $h_{j}$’s are holomorphic on $V.$
 Proof.
@@
 Lemma 3.4.
-Assume that there is a subsequence $r_{j_{k}}$ of $r_{j}$ such that for each $k$ there exists a point ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j_{k}})}$ with $|p_{1}(a_{k},b_{k})|=r_{j_{k}}$ and $|p_{2}(a_{k},b_{k})|<r_{j_{k}}-\varepsilon_{j_{k}}.$ Then there exists a analytic set normal-$\triangle$ and all $h_{j}$’s are holomorphic on .normal-$\triangle$.△ .
+Assume that there is a subsequence $r_{j_{k}}$ of $r_{j}$ such that for each $k$ there exists a point ${\sf Gr}_{h}(a_{k},b_{k})\in\widehat{{\sf Gr}_{h}(\mathfrak{Q}_{j_{k}})}$ with $|p_{1}(a_{k},b_{k})|=r_{j_{k}}$ and $|p_{2}(a_{k},b_{k})|<r_{j_{k}}-\varepsilon_{j_{k}}.$ Then there exists a analytic set $\triangle$ and all $h_{j}$’s are holomorphic on $\triangle$.△ .
 Proof.
@@
 Note that $\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k})\}$ is equivalent to $\mathbb{C}$ (by Result 2.8), and $u_{j}$’s are harmonic on $\triangle_{k}.$ Hence $u$ is constant on some component of $\triangle_{k},$ i.e.
-all $h_{j}$’s are holomorphic on that component of $\triangle_{k}.$ We will denote this component again by $\triangle_{k}.$ We now claim that ${\sf Gr}_{h}(\triangle_{k})\subset\widehat{{\sf Gr}_{h}(\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\})}.$ To prove this claim, let $(\xi,h(\xi))\in{\sf Gr}_{h}(\triangle_{k})$ for some $\xi\in\triangle_{k}.$ Then for any polynomial $P(z,w)$ in $\mathbb{C}^{2},$
+all $h_{j}$’s are holomorphic on that component of $\triangle_{k}.$ We will denote this component again by $\triangle_{k}.$
+We now claim that ${\sf Gr}_{h}(\triangle_{k})\subset\widehat{{\sf Gr}_{h}(\{p_{1}(z)=e^{i\theta_{0}}p_{1}(a_{k},b_{k}),|p_{2}(z)|=r_{j_{k}}\})}.$ To prove this claim, let $(\xi,h(\xi))\in{\sf Gr}_{h}(\triangle_{k})$ for some $\xi\in\triangle_{k}.$ Then for any polynomial $P(z,w)$ in $\mathbb{C}^{2},$
 $|P(\xi,h(\xi))|$ $\leq\sup_{{\sf Gr}_{h}(\triangle_{k})}|P(z,w)|$
@@
 $\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}\cap F^{-1}(0)$ $=\widehat{({\sf Gr}_{u}(\mathfrak{Q}_{j})\cap F^{-1}\{0\}})=\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})}.$
-Since ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}$ and ${\sf Gr}_{u}(\alpha_{0})\in F^{-1}\{0\},$ hence ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})}.$ Now,
+Since ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}$ and ${\sf Gr}_{u}(\alpha_{0})\in F^{-1}\{0\},$ hence ${\sf Gr}_{u}(\alpha_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}^{{\bf c}}_{j})}.$
+Now,
 $\mathfrak{Q}^{{\bf c}}_{j}\subset\widetilde{L_{{\bf c}}}:=\{z\in\overline{\mathfrak{D}^{r_{j}}_{2}}:u(z)={\bf c}\}$
@@
 Therefore, $\alpha_{0}\in\widehat{\mathfrak{Q}^{\bf c}_{j}}=\mathfrak{Q}^{\bf c}_{j},$ but we have taken $\alpha_{0}$ from $\mathfrak{D}^{r_{j}}_{2}.$ This is a contradiction.
-Therefore, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains at least one point $z_{0}\in L_{c_{1}}\setminus W.$ We claim that ${\sf Gr}_{u}(z_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}.$ To prove this claim, let $H(z,w)$ be a polynomial in $\mathbb{C}^{2+N}.$ Since $(z_{0},u(z_{0}))\in{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})$, then
+Therefore, $\widehat{\mathfrak{Q}^{\bf c}_{j}}$ contains at least one point $z_{0}\in L_{c_{1}}\setminus W.$
+We claim that ${\sf Gr}_{u}(z_{0})\in\widehat{{\sf Gr}_{u}(\mathfrak{Q}_{j})}.$ To prove this claim, let $H(z,w)$ be a polynomial in $\mathbb{C}^{2+N}.$ Since $(z_{0},u(z_{0}))\in{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})$, then
 $|H(z_{0},u(z_{0}))|$ $\leq\sup_{{\sf Gr}_{u}(\widehat{\mathfrak{Q}^{\bf c}_{j}})}|H(z,w)|$
@@
 $\subset(L_{\bf c}\cap\bar{U}\setminus D(z_{0}))\cup(\mathfrak{L}_{z_{0}}\cap\partial U).$
-Case I If $s\in\mathfrak{L}_{z_{0}}\cap\partial U$ and since $s\in X,$ therefore $s\in X_{z_{0}}=X\cap\partial D(z_{0}),$ then we have $|\gamma(s)|<1.$ Also, since $\Phi_{1}(s)\in i\mathbb{R},$ there exists $y\in\mathbb{R}$ such that $\Phi_{1}(s)=iy.$ Therefore, $(\xi\circ\Phi_{1}(s))=e^{-y^{2}}<1.$ This implies $|(\xi\circ\Phi_{1}(s))^{d}|\leq 1.$ This contradicts that $z_{0}\in\widehat{X}.$ Case II If $s\in(L_{\bf c}\cap\bar{U}\setminus D(z_{0}))$ and $s\notin\mathfrak{L}_{z_{0}}\cap\partial U$ implies $s\notin\mathfrak{L}_{z_{0}},$ by (3.10), $|\xi\circ\Phi_{1}(s))|<1.$ Then for sufficiently large $d$, we have $|\gamma(s)|\times|\xi\circ\Phi_{1}(s))|^{d}<1.$ Again this is a contradiction to the fact that $z_{0}\in\widehat{X}.$ Therefore, $X_{z_{0}}=\partial D(z_{0})\subset X.$ Hence,
+Case I If $s\in\mathfrak{L}_{z_{0}}\cap\partial U$ and since $s\in X,$ therefore $s\in X_{z_{0}}=X\cap\partial D(z_{0}),$ then we have $|\gamma(s)|<1.$ Also, since $\Phi_{1}(s)\in i\mathbb{R},$ there exists $y\in\mathbb{R}$ such that $\Phi_{1}(s)=iy.$ Therefore, $(\xi\circ\Phi_{1}(s))=e^{-y^{2}}<1.$ This implies $|(\xi\circ\Phi_{1}(s))^{d}|\leq 1.$ This contradicts that $z_{0}\in\widehat{X}.$
+Case II If $s\in(L_{\bf c}\cap\bar{U}\setminus D(z_{0}))$ and $s\notin\mathfrak{L}_{z_{0}}\cap\partial U$ implies $s\notin\mathfrak{L}_{z_{0}},$ by (3.10), $|\xi\circ\Phi_{1}(s))|<1.$ Then for sufficiently large $d$, we have $|\gamma(s)|\times|\xi\circ\Phi_{1}(s))|^{d}<1.$ Again this is a contradiction to the fact that $z_{0}\in\widehat{X}.$ Therefore, $X_{z_{0}}=\partial D(z_{0})\subset X.$ Hence,
 $D(z_{0})\subset\widehat{\partial D(z_{0})}\subset\widehat{X}\subset\widehat{\mathfrak{Q}^{\bf c}_{j}}\subset\widetilde{L_{\bf c}}.$ (3.11)
@@
 $\widetilde{V}:=\{z\in\mathfrak{D}_{2}:\bar{\partial}{{h_{j_{1}}}}\wedge\bar{\partial}{{h_{j_{2}}}}(z)=0,\forall~{}{j_{1},j_{2}},1\leq j_{1},j_{2}\leq N\}.$
-Since, by assumption in the beginning of the proof, ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, by (3.4), there exist $\alpha_{0}\in\widetilde{V}$ such that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Case I: Assume that $\widetilde{V}\not=\mathfrak{D}_{2}.$ Then Lemma 3.2 gives an irreducible component $V$ of $\widetilde{V}$ such that all $h_{j}$’s are holomorphic on $V.$ Observe that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ The proof of the theorem is complete in this case in view of Lemma 2.22.
+Since, by assumption in the beginning of the proof, ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, by (3.4), there exist $\alpha_{0}\in\widetilde{V}$ such that ${\sf Gr}_{h}(\alpha_{0})\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$
+Case I: Assume that $\widetilde{V}\not=\mathfrak{D}_{2}.$ Then Lemma 3.2 gives an irreducible component $V$ of $\widetilde{V}$ such that all $h_{j}$’s are holomorphic on $V.$ Observe that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ The proof of the theorem is complete in this case in view of Lemma 2.22.
 Case II: Now we consider the case that $\widetilde{V}=\mathfrak{D}_{2}.$ Without loss of generality, we assume that $h_{1}$ is non-holomorphic.
@@
 If #$\#CC(X)=\infty,$ then there is nothing to prove.
-In view Result 2.11, it is not possible that #$\#CC(X)<\infty,$ but #$\#CC(\widehat{X})=\infty.$ In fact, assume that $\widehat{X}=\cup^{\infty}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ This implies $X=X\cap\widehat{X}=\cup^{\infty}_{j=1}X_{j}\cap X.$ Result 2.11 says that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ Since each $X_{j}\cap X$ is disjoint, and #$\#CC(X)<\infty,$ therefore there exists $j_{0}\in\mathbb{N}$ such that $X_{j}\cap X=\emptyset~{}\forall j>j_{0}.$ Hence $X_{j}=\widehat{X_{j}\cap X}=\emptyset~{}\forall j>j_{0}.$ This is a contradiction to the fact that #$\#CC(\widehat{X})=\infty.$ Next, we assume that #$\#CC(X)=m$ and #$\#CC(\widehat{X})=l.$ Assume that $\widehat{X}=\cup^{l}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ Then $X=\widehat{X}\cap X=\cup^{l}_{j=1}X_{j}\cap X.$ By using Result 2.11, we obtain that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ If each $X_{j}\cap X$ is connected, then $l=m.$ Otherwise $l<m.$
+In view Result 2.11, it is not possible that #$\#CC(X)<\infty,$ but #$\#CC(\widehat{X})=\infty.$ In fact, assume that $\widehat{X}=\cup^{\infty}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ This implies $X=X\cap\widehat{X}=\cup^{\infty}_{j=1}X_{j}\cap X.$ Result 2.11 says that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ Since each $X_{j}\cap X$ is disjoint, and #$\#CC(X)<\infty,$ therefore there exists $j_{0}\in\mathbb{N}$ such that $X_{j}\cap X=\emptyset~{}\forall j>j_{0}.$ Hence $X_{j}=\widehat{X_{j}\cap X}=\emptyset~{}\forall j>j_{0}.$ This is a contradiction to the fact that #$\#CC(\widehat{X})=\infty.$
+Next, we assume that #$\#CC(X)=m$ and #$\#CC(\widehat{X})=l.$ Assume that $\widehat{X}=\cup^{l}_{j=1}X_{j},$ where each $X_{j}$ is disjoint connected component of $\widehat{X}.$ Then $X=\widehat{X}\cap X=\cup^{l}_{j=1}X_{j}\cap X.$ By using Result 2.11, we obtain that $X_{j}=\widehat{X_{j}\cap X}~{}\forall j\geq 1.$ If each $X_{j}\cap X$ is connected, then $l=m.$ Otherwise $l<m.$
 We write
@@
 $\triangle(z)=\begin{vmatrix}\frac{\partial P(\Psi)}{\partial{z_{1}}}&\frac{\partial P(\Psi)}{\partial z_{2}}\\[6.45831pt] \frac{\partial h(\Psi)}{\partial{{z_{1}}}}&\frac{\partial h(\Psi)}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}\times\begin{vmatrix}\frac{\partial p_{1}}{\partial{{z_{1}}}}&\frac{\partial p_{1}}{\partial{z_{2}}}\\[6.45831pt] \frac{\partial p_{2}}{\partial{{z_{1}}}}&\frac{\partial p_{2}}{\partial{z_{2}}}\\[6.45831pt] \end{vmatrix}_{z}=\frac{1}{p^{m+1}_{1}p^{n+1}_{2}}\Pi^{l}_{j=1}q_{j}(z),$
-for some $l\in\mathbb{N},$ where each $q_{j}$ is an irreducible polynomial in $\mathbb{C}^{2}.$ We define the corresponding irreducible algebraic variety $Z_{j}:=\{z\in\mathbb{C}^{2}:q_{j}(z)=0\}.$ We assume $\triangle({z})\not\equiv 0$ on $X.$ Therefore, each $q_{j}$ is a non-zero holomorphic polynomial in $\mathbb{C}^{2}.$ We set
+for some $l\in\mathbb{N},$ where each $q_{j}$ is an irreducible polynomial in $\mathbb{C}^{2}.$ We define the corresponding irreducible algebraic variety $Z_{j}:=\{z\in\mathbb{C}^{2}:q_{j}(z)=0\}.$ We assume $\triangle({z})\not\equiv 0$ on $X.$ Therefore, each $q_{j}$ is a non-zero holomorphic polynomial in $\mathbb{C}^{2}.$
+We set
 $\Sigma:=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\triangle{(z)}=0\right\}.$ (4.1)
@@
 $|g(z_{1},z_{2},\xi)|<\frac{1}{2}\text{ on }{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}}).$
-Assume that $\sup_{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}|g(z_{1},z_{2},\xi)|=|g(s,\overline{P\circ\Psi}(s))|$ for some $s=(s_{1},s_{2})\in\Gamma_{{\mathfrak{D}_{2}}}.$ Case I: $s\in\Gamma_{{\mathfrak{D}_{2}}}\setminus(U\cup W),$ then there is nothing to prove.
-Case II: $s\in U$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$ Case III: $s\in W$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$ Combining all the three cases, we can conclude that ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$ Next, we assume that $a_{0}\in(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}).$ Since $(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})$ is finite and $\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}$ does not containing any isolated point, therefore, in this case also ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
+Assume that $\sup_{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{{\mathfrak{D}_{2}}})}|g(z_{1},z_{2},\xi)|=|g(s,\overline{P\circ\Psi}(s))|$ for some $s=(s_{1},s_{2})\in\Gamma_{{\mathfrak{D}_{2}}}.$
+Case I: $s\in\Gamma_{{\mathfrak{D}_{2}}}\setminus(U\cup W),$ then there is nothing to prove.
+Case II: $s\in U$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$
+Case III: $s\in W$ and $s\in\Gamma_{{\mathfrak{D}_{2}}}.$ Then $|g_{1}(s,\overline{P\circ\Psi}(s))|\leq 1$ but $|p_{0}(s)f_{2}(s)|<\frac{1}{2}.$ This implies $|g(s,\overline{P\circ\Psi}(s))|<\frac{1}{2}.$
+Combining all the three cases, we can conclude that ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
+Next, we assume that $a_{0}\in(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}}).$ Since $(Z_{j_{0}}\cap\overline{\mathfrak{D}}_{2}\cap A)\setminus(L\cup\Gamma_{{\mathfrak{D}_{2}}})$ is finite and $\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}$ does not containing any isolated point, therefore, in this case also ${\sf Gr}_{\overline{(P\circ\Psi)}}(a_{0})\notin\widehat{{\sf Gr}_{\overline{(P\circ\Psi)}}(\Gamma_{{\mathfrak{D}_{2}}})}.$
 In view of Lemma 4.10 and Lemma 4.9, we get that
@@
 $[z_{1},z_{2},f_{1},\cdots,f_{N};\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}),$
-or there exists a non-trivial algebraic variety $V\subset\mathbb{C}^{2}$ with $V\cap\partial\mathfrak{D}_{2}\subset\Gamma_{\mathfrak{D}_{2}},$ and the pluriharmonic extensions of the $f_{j}$ are holomorphic on $V.$ Explanation: We claim that
+or there exists a non-trivial algebraic variety $V\subset\mathbb{C}^{2}$ with $V\cap\partial\mathfrak{D}_{2}\subset\Gamma_{\mathfrak{D}_{2}},$ and the pluriharmonic extensions of the $f_{j}$ are holomorphic on $V.$
+Explanation: We claim that
   1. (i)
@@
 […]
+Proof of (iii): It is easy to see that $\overline{\mathfrak{D}}_{2}$ is path-connected.
+To see this, let $(a,b)\in\overline{\mathfrak{D}}_{2}.$ Then $\eta(t):=(t^{l}a,tb):[0,1]\to\overline{\mathfrak{D}}_{2}$ is a path between $(a,b)$ and the origin.
+Therefore, there is always a path between the origin and the arbitrary point of $\overline{\mathfrak{D}}_{2}.$ Hence $\overline{\mathfrak{D}}_{2}$ is path connected.
+Now we show that $\overline{\mathfrak{D}}_{2}$ is simply-connected.
+For this, let $\gamma=(\gamma_{1},\gamma_{2}):[0,1]\to\overline{\mathfrak{D}}_{2}$ be a loop in $\overline{\mathfrak{D}}_{2}$ based at $(0,0).$ We consider the function $H(t,x):[0,1]\times[0,1]\to\overline{\mathfrak{D}}_{2}$ by
 $H(t,x)=\left(t^{l}\gamma_{1}(x),t\gamma_{2}(x)\right).$
@@
 Hence polynomial approximation does not holds.
-Explanation: Following the notation in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{1}{z_{2}}.$ Therefore,
+Explanation: Following the notations in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{1}{z_{2}}.$ Therefore,
 $X$ =$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{(P\circ\Psi)(z)}=(h\circ\Psi)(z)\right\}$
@@
 $=\frac{1}{p^{2}_{1}p^{2}_{2}}(p_{1}+p_{2})(p_{1}-p_{2})(z_{2}-z_{1}).$
-We define $q_{1}:=p_{1}-p_{2}=z_{1}+z_{2}-z_{1}z_{2},~{}~{}q_{2}:=p_{1}+p_{2}=z_{1}+z_{2}+z_{1}z_{2},~{}~{}q_{3}=z_{2}-z_{1}$ and $Z_{j}=\{q_{j}=0\}.$ Therefore, Clearly, $Z_{j}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ only for ${j}=2.$ It is easy to see that $\widehat{Q}_{2}=\widehat{Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+z_{2}+z_{1}z_{2}=0\}.$ Indeed, we write $\mathcal{K}=\{(z_{1},z_{2})\in\mathbb{C}^{2}:z_{1}+z_{2}=0,|z_{1}|=1,|z_{2}|=1\}.$ Clearly, $\Psi^{-1}(\mathcal{K})=Q_{2}=Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ We now claim that
+We define $q_{1}:=p_{1}-p_{2}=z_{1}+z_{2}-z_{1}z_{2},~{}~{}q_{2}:=p_{1}+p_{2}=z_{1}+z_{2}+z_{1}z_{2},~{}~{}q_{3}=z_{2}-z_{1}$ and $Z_{j}=\{q_{j}=0\}.$ Therefore,
+Clearly, $Z_{j}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ only for ${j}=2.$ It is easy to see that $\widehat{Q}_{2}=\widehat{Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+z_{2}+z_{1}z_{2}=0\}.$ Indeed, we write $\mathcal{K}=\{(z_{1},z_{2})\in\mathbb{C}^{2}:z_{1}+z_{2}=0,|z_{1}|=1,|z_{2}|=1\}.$ Clearly, $\Psi^{-1}(\mathcal{K})=Q_{2}=Z_{2}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ We now claim that
 $\Psi^{-1}(\widehat{\mathcal{K}})=\widehat{Q_{2}}.$
@@
 Thus $\widehat{\Psi(Q_{2})}=\widehat{\Psi(\widehat{Q_{2}})}=\Psi(\widehat{Q_{2}}).$ This implies $\widehat{Q_{2}}=\Psi^{-1}(\widehat{\mathcal{K}}),$ i.e.
-$\widehat{Q_{2}}=\overline{\mathfrak{D}}_{2}\cap Z_{2}.$ Therefore, by Theorem 4.3, we get that
+$\widehat{Q_{2}}=\overline{\mathfrak{D}}_{2}\cap Z_{2}.$
+Therefore, by Theorem 4.3, we get that
 $\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})\cup\{(z_{1},z_{2},0):(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2},z_{1}+z_{2}+z_{1}z_{2}=0\}.$
@@
 $=\frac{1}{p^{2}_{1}p^{2}_{2}}(p_{1}+p_{2})(p_{1}-p_{2})(-6z_{2}).$
-We define $q_{1}:=p_{1}-p_{2}=z_{1}+2z^{2}_{2},~{}~{}q_{2}:=p_{1}+p_{2}=3z_{1},~{}~{}q_{3}=-6z_{2}$ and $Z_{j}=\{q_{j}=0\}.$ Clearly, $Q_{3}=Z_{3}\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset$ and it is easy to show that if $z\in Z_{2}\cap X,$ then $z\in\Gamma_{{\mathfrak{D}_{2}}},$ this is not possible because $X\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset.$ Now its remaining to verify for the set $Q_{1}=Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ It is easy to see that $Z_{1}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ and $\widehat{Q}_{1}=\widehat{Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+2z^{2}_{2}=0\}.$ Therefore, by Theorem 4.3, we get that
+We define $q_{1}:=p_{1}-p_{2}=z_{1}+2z^{2}_{2},~{}~{}q_{2}:=p_{1}+p_{2}=3z_{1},~{}~{}q_{3}=-6z_{2}$ and $Z_{j}=\{q_{j}=0\}.$
+Clearly, $Q_{3}=Z_{3}\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset$ and it is easy to show that if $z\in Z_{2}\cap X,$ then $z\in\Gamma_{{\mathfrak{D}_{2}}},$ this is not possible because $X\cap\Gamma_{{\mathfrak{D}_{2}}}=\emptyset.$ Now its remaining to verify for the set $Q_{1}=Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}.$ It is easy to see that $Z_{1}\cap\overline{\mathfrak{D}}_{2}\setminus(\Gamma_{{\mathfrak{D}_{2}}}\cup L)\subset X$ and $\widehat{Q}_{1}=\widehat{Z_{1}\cap\Gamma_{{\mathfrak{D}_{2}}}}=\{(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2}:z_{1}+2z^{2}_{2}=0\}.$ Therefore, by Theorem 4.3, we get that
 $\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})\cup\{(z_{1},z_{2},0):(z_{1},z_{2})\in\overline{\mathfrak{D}}_{2},z_{1}+2z^{2}_{2}=0\}.$
@@
 Let $p_{1}(z_{1},z_{2})=2z_{1}+z^{2}_{2},~{}~{}p_{2}(z_{1},z_{2})=z_{1}-z^{2}_{2},~{}~{}P(z_{1},z_{2})=z_{1}+2z_{2}$ and $\Psi(z_{1},z_{2})=(p_{1}(z_{1},z_{2}),p_{2}(z_{1},z_{2})).$ Then the graph of $\overline{P\circ\Psi}$ over $\Gamma_{{\mathfrak{D}_{2}}}$ is polynomially convex.
-Explanation: Following the notation in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{2}{z_{2}}.$ Therefore,
+Explanation: Following the notations in Theorem 4.3, let $\Psi(z)=(p_{1}(z),p_{2}(z)),$ $h(z)=\frac{1}{z_{1}}+\frac{2}{z_{2}}.$ Therefore,
 $X$ =$=\left\{z\in\overline{\mathfrak{D}_{2}}\setminus(L\cup\Gamma_{\mathfrak{D}_{2}}):\overline{(P\circ\Psi)(z)}=(h\circ\Psi)(z)\right\}$
@@
 $\widehat{{\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}})}={\sf Gr}_{\overline{P\circ\Psi}}(\Gamma_{\mathfrak{D}_{2}}).$
-Therefore, by using Theorem 2.24, we get $[z_{1},z_{2},\overline{(P\circ\Psi)}(z);\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$ Acknowledgments.
+Therefore, by using Theorem 2.24, we get $[z_{1},z_{2},\overline{(P\circ\Psi)}(z);\Gamma_{\mathfrak{D}_{2}}]=\mathcal{C}(\Gamma_{\mathfrak{D}_{2}}).$
+Acknowledgments.
 The first named author was partially supported by a MATRICS Research Grant (MTR/2017/000974) of SERB, Dept.
 of Science and Technology, Govt.
-of India, for the beginning of this work and is supported partially by a Core Research Grant (CRG/2022/003560) of SERB, Dept.
-of Science and Technology, Govt.
+of India, for the beginning of this work and is supported partially by a Core Research Grant (CRG/2022/003560) of ANRF (formerly SERB), Govt.
 of India, for the later part of the work.
@@
 of India, for the early version of this work, and an IoE-IISc Postdoctoral Fellowship (R(HR)(IOE-PDF)(MA)(GMM)-114) for the latest version of this work.
-The second author would also like to thank Amar Deep Sarkar for the fruitful discussions.
+The second named author would also like to thank Amar Deep Sarkar for the fruitful discussions.